### v1

# Waring and Waring–Goldbach subbases with prescribed representation function

Abstract.
We study representations of integers as sums of $h$ elements from subsets of $k$-th powers $\mathbb{N}^{k}$ and $k$-th powers of primes $\mathbb{P}^{k}$ for $k\geq 1$. Extending work by Vu, Wooley, and others, we show that for $h\geq h_{k}=O(8^{k}k^{2})$, if $F$ is a regularly varying function satisfying $\lim_{n\to\infty}F(n)/\log n=\infty$ and certain growth conditions, then there exists $A\subseteq\mathbb{N}^{k}$ such that $r_{A,h}(n)\sim\mathfrak{S}_{k,h}(n)F(n)$, where $\mathfrak{S}_{k,h}(n)$ is the singular series from Waring’s problem.
For $h\geq 2^{k}+1$, we show the existence of $A\subseteq\mathbb{P}^{k}$ with $r_{A,h}(n)\sim\mathfrak{S}^{*}_{k,h}(n)cn^{\kappa}$ for $10<\kappa<h/k-$ 1 and $c>0$, where $\mathfrak{S}^{*}_{k,h}(n)$ is the singular series from Waring–Goldbach’s problem. Additionally, for $10\leq\kappa\leq h/k-$ 1 and functions $\psi$ satisfying $\psi(x)\asymp_{\lambda}\psi(x^{\lambda})$ for every $\lambda>0$, if $\log x\ll x^{\kappa}\psi(x)\ll x^{h/k-1}/(\log x)^{h}$, then there exists $A\subseteq\mathbb{P}^{k}$ with $r_{A,h}(n)\asymp n^{\kappa}\psi(n)$ for $n$ satisfying certain congruence conditions, producing thin subbases of prime powers when $\kappa=0$, $\psi=\log$.
Key words and phrases:
probabilistic method, regular variation, Waring’s problem, Waring–Goldbach
2020 Mathematics Subject Classification:
11P05, 11P32, 11B13, 11B34, 05D40
1. Introduction
Let $A\subseteq\mathbb{Z}_{\geq 0}$ and $h\geq 2$ be an integer. Define the representation function
$r_{A,h}(n):=\{(x_{1},\ldots,x_{h})\in A^{h}\mid x_{1}+\cdots+x_{h}=n\},$
which counts the number of ways an integer $n$ can be expressed as a sum of $h$ elements from $A$. In 1956, inspired by a problem of Sidon, Erdős [3] used probabilistic methods to show that there exists a subset $A\subseteq\mathbb{N}$ such that $r_{A,2}(n)\asymp\log n$. Later work by Erdős [4, Corollary 3] established that for certain functions $F$ satisfying $\lim_{x\to\infty}F(x)/\log x=\infty$ (e.g., $F(x)=\log n\log\log n$), there exist subsets $A\subseteq\mathbb{N}$ such that $r_{A,2}(n)\sim F(n)$.
In 1990, Erdős and Tetali [5] extended this result for any $h\geq 2$, showing the existence of subsets $A\subseteq\mathbb{N}$ such that $r_{A,h}(n)\asymp\log n$. In our earlier work [14], we extended this line of research by studying subsets $A$ whose representation functions grow according to regularly varying functions. Specifically, we proved that for any regularly varying function $F$ such that
$\lim_{x\to\infty}\frac{F(x)}{\log x}=\infty\quad\text{ and }\quad F(x)\leq(1+o(1))\frac{x^{h-1}}{(h-1)!},$
there exists $A\subseteq\mathbb{N}$ such that $r_{A,h}(n)\sim F(n)$.
A real-valued function $F:\mathbb{R}_{>0}\to\mathbb{R}_{>0}$ is called regularly varying if it is measurable and satisfies $\lim_{x\to\infty}F(\lambda x)/F(x)$ for every $\lambda>0$. Regularly varying functions are of the form
$F(x)=x^{\kappa}\psi(x),$
where $\kappa\in\mathbb{R}$ and $\psi(x)$ is slowly varying, meaning $\lim_{x\to\infty}\psi(\lambda x)/\psi(x)=1$ for any $\lambda>0$. Slowly varying functions satisfy $\psi(x)=x^{o(1)}$ (see Bingham–Goldie–Teugels [1]). These functions provide a broad and flexible framework for describing growth conditions, encompassing a variety of natural cases.
In this paper, we expand upon these results by considering subsets of $k$-th powers $\mathbb{N}^{k}:=\{n^{k}\leavevmode\nobreak\ |\leavevmode\nobreak\ n\in\mathbb{N}\}$ and $k$-th powers of primes $\mathbb{P}^{k}:=\{p^{k}\leavevmode\nobreak\ |\leavevmode\nobreak\ p\text{ is prime}\}$.
1.1. Waring subbases
In 2000, Vu [17] showed that for each $k\geq 2$, there exists an integer $h_{k}\ll 8^{k}k^{2}$ such that for all $h\geq h_{k}$, there exists a subset $A\subseteq\mathbb{N}^{k}$ for which $r_{A,h}(n)\asymp\log n$. Later, Wooley [19] refined this bound, proving that
$h_{k}\leq k\left(\log k+\log\log k+2+O\left(\frac{\log\log k}{\log k}\right)\right).$
More recently, drawing from the work of Brüdern–Wooley [2], Pliego [13] further improved the bound to
$h_{k}\leq\lceil k(\log k+4.20032)\rceil,$
showing that bounds for $h_{k}$ essentially align with those for the asymptotic order of $\mathbb{N}^{k}$, as derived using the circle method. Pliego also established that for certain functions $\psi(x)=x^{o(1)}$ with $\lim_{x\to\infty}\psi(x)/\log x=\infty$ (e.g., $\psi(x)=\log n\log\log n$), there exist $A\subseteq\mathbb{N}^{k}$ such that
$r_{A,h}(n)\sim\mathfrak{S}_{k,h}(n)\psi(n),$
where $\mathfrak{S}_{k,h}(n)$ denotes the singular series associated with Waring’s problem, defined by
$S(a,q):=\sum_{r=1}^{q}e\bigg{(}\frac{ar^{k}}{q}\bigg{)},\qquad\mathfrak{S}_{k,h}(n):=\sum_{q\geq 1}\sum_{\begin{subarray}{c}a=1\\ (a,q)=1\end{subarray}}^{q}\frac{S(a,q)^{h}}{q^{h}}e\bigg{(}-\frac{na}{q}\bigg{)}.$ (1.1)
For $h\geq 4k$, this series converges absolutely, and $\mathfrak{S}_{k,h}(n)\asymp 1$ (cf. Vaughan [15, Theorems 4.3, 4.6]).
For values of $h$ as large as those in Vu’s work, we show the existence of subbases of $\mathbb{N}^{k}$ with prescribed asymptotics:
Theorem 1.1.
Let $k\geq 2$ be an integer.
  1. (i)
    There exists $h_{k}\leq k^{2}(\log k+\log\log k+O(1))$ such that the following holds: Let $h\geq h_{k}$, and $10<\kappa<h/k-$ 1. Then, for every $c>0$, there exists a subset $A\subseteq\mathbb{P}^{k}$ such that
    $r_{A,h}(n)\sim\mathfrak{S}_{k,h}(n)\,cn^{\kappa}.$
    The same holds for $\kappa=h/k-1$ but with $0<c\leq\frac{\Gamma(1+1/k)^{h}}{\Gamma(h/k)}$.
  2. (ii)
    There exists $h_{k}^{\prime}\ll 8^{k}k^{2}$ such that the following holds: Let $h\geq h^{\prime}_{k}$, and $F$ be a regularly varying function satisfying
    $\lim_{x\to\infty}\frac{F(x)}{\log x}=\infty,\qquad F(x)\leq(1+o(1))\frac{\Gamma(1+1/k)^{h}}{\Gamma(h/k)}x^{h/k-1},$
    such that $F(x)/x^{h/k-1}$ is asymptotic to a non-increasing function. Then, there exists a subset $A\subseteq\mathbb{N}^{k}$ such that
    $r_{A,h}(n)\sim\mathfrak{S}_{k,h}(n)F(n).$
In both cases, $\mathfrak{S}_{k,h}(n)$ is the singular series for Waring’s problem (1.1).
Remark.
In Theorem 1.1 (ii), if $F(x)\ll x^{h/k-1-\varepsilon}$ for some $\varepsilon>0$, then $F(x)/x^{h/k-1}$ is necessarily asymptotic to a non-increasing function, since $F(x)/x^{h/k-1}\sim\sup_{y\geq x}F(y)/y^{h/k-1}$ (cf. BGT [1, Theorem 1.5.3]).
1.2. Waring–Goldbach subbases
Let $k\geq 1$. Given a prime $p$, let $\theta=\theta(k,p)$ be the unique integer such that $p^{\theta}\mid k$ but $p^{\theta+1}\nmid k$. Define
$\gamma=\gamma(k,p):=\begin{cases}\theta+2&\text{ if }p=2,\,2\mid k,\\ \theta+1&\text{ otherwise},\end{cases}\qquad K(k):=\prod_{(p-1)\mid k}p^{\gamma}.$ (1.2)
If a prime $q$ is coprime to $K(k)$, then $q^{k}\equiv 1\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ p^{\gamma})$ whenever $(p-1)\mid k$, since $\varphi(p^{\gamma})=p^{\theta}(p-1)\mid k$ for $p$ odd (and $\varphi(2^{\gamma})/2\mid k$), where $\varphi(n):=|\{m\leq n\leavevmode\nobreak\ |\leavevmode\nobreak\ (m,n)=1\}|$ is Euler’s totient function. So by the Chinese remainder theorem, if $n$ is a sum of $h$ $k$-th powers of primes greater than $k+1$, then $n\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$. Note that if $k$ is odd, then $K(k)=2$.
$k$ 1 2 3 4 5 6 7 8 9 10
$K(k)$ 2 24 2 240 2 504 2 480 2 264
Table 1. First few values of $K(k)$.
The singular series associated to Waring–Goldbach’s problem is defined as
$\mathfrak{S}^{*}_{k,h}(n):=\sum_{q\geq 1}\sum_{\begin{subarray}{c}a=1\\ (a,q)=1\end{subarray}}^{q}\frac{S(q,a)^{h}}{\varphi(q)^{h}}e\bigg{(}-\frac{na}{q}\bigg{)},$ (1.3)
where $S(q,a)$ is as in (1.1). For $h\geq 2^{k}+1$, we have $\mathfrak{S}^{*}_{k,h}(n)\asymp 1$ for $n\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$ (see Lemma 5.4). In order to obtain thin subbases of prime powers, we first prove a result similar to Wooley [20, Theorem 1.1] for $\mathbb{P}^{k}$.
Theorem 1.2.
Let $k\geq 1$ be an integer, and $h\geq 2^{k}+1$. For every $\omega\geq 1/h$, we have
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{P}^{k}\\ x_{1}+\cdots+x_{h}=n\end{subarray}}(x_{1}\cdots x_{h})^{\omega-\frac{1}{k}}(\log x_{1}\cdots\log x_{h})=\mathfrak{S}^{*}_{k,h}(n)\,\frac{\Gamma(\omega)^{h}}{\Gamma(h\omega)}\,n^{h\omega-1}+O_{R}\bigg{(}\frac{n^{h\omega-1}}{(\log n)^{R}}\bigg{)}$
for every $R>1$.
Note that the implied constant also depends on $h$, $k$, but we omit this dependency. With this, we will show the following:
Theorem 1.3.
Let $k\geq 1$ be an integer, $h\geq 2^{k}+1$, and suppose that $n\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$.
  1. (i)
    For every $10<\kappa<h/k-$ 1 and $c>0$, there exists a subset $A\subseteq\mathbb{P}^{k}$ such that
    $r_{A,h}(n)\sim\mathfrak{S}^{*}_{k,h}(n)\,cn^{\kappa},$
    where $\mathfrak{S}^{*}_{k,h}(n)$ is the singular series for Waring–Goldbach’s problem (1.3).
  2. (ii)
    Let $10\leq\kappa\leq h/k-$ 1, $\psi$ be a measurable positive real function satisfying $\psi(x)\asymp_{\lambda}\psi(x^{\lambda})$ for every $\lambda>0$, and write $F(x)=x^{\kappa}\psi(x)$. If $\log x\ll F(x)\ll x^{h/k-1}/(\log x)^{h}$, then there exists a subset $A\subseteq\mathbb{P}^{k}$ such that
    $r_{A,h}(n)\asymp F(n).$
[…]
Remark.
[…]
Assuming the asymptotic version of Goldbach’s conjecture, Granville [6, Theorem 2] showed that there is a subset $A\subseteq\mathbb{P}$ with $|A\cap[1,x]|\asymp(x\log x)^{1/2}$ such that every large even integer is the sum of two elements of $A$. Our result, however, does not apply to this case.
1.3. General setup
Theorems 1.1 and 1.3 will be derived from the more general Theorem 1.4 below. Let $B\subseteq\mathbb{Z}_{\geq 0}$ be a subset, and write $B(x):=|B\cap[1,x]|$. Let $F$ be a regularly varying function. We make the following assumptions about $B$ and $F$:
  1. (i)
    (Regular variation) There is $\beta=\beta(B)>0$ such that, for every fixed $\lambda\geq 1$,
    $B(\lambda x)\sim\lambda^{\beta}B(x).$
  2. (ii)
    (Low additive energy) There exists $H_{\textnormal{(ii)}}\in\mathbb{Z}_{\geq 1}$ such that, for every $h\geq H_{\textnormal{(ii)}}$,
    $\sum_{n\leq x}r_{B,h}(n)^{2}\ll\frac{B(x)^{2h+o(1)}}{x}.$
  3. (iii)
    (Counting solutions with weights) There exists $H_{\textnormal{(iii)}}\in\mathbb{Z}_{\geq 1}$ such that the following holds: Let $h\geq H_{\textnormal{(iii)}}$, and suppose $F(n)\leq(1+o(1))r_{B,h}(n)$ is such that $f(x):=(xF(x))^{1/h}=x^{\omega}\phi(x)$ satisfies $1/h\leq\omega\leq\beta$, with $f(x)/B(x)$ asymptotic to a non-increasing function. Then:
    $\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in B\\ x_{1}+\cdots+x_{h}=n\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots\frac{f(x_{h})}{B(x_{h})}\sim\mathfrak{S}_{B,h}(n)\,C_{B,h,f}\,\frac{f(n)^{h}}{n},$
    where $C_{B,h,f}\in\mathbb{R}_{>0}$ is a constant, and $\mathfrak{S}_{B,h}(n)$ is some function of $n$ satisfying
    $\mathfrak{S}_{B,h}(n)\asymp 1\text{ for }n\in\mathscr{S},\text{ for some subset }\mathscr{S}\subseteq\mathbb{Z}_{\geq 0}.$
Condition (ii) asks for $hB$ to have almost as low additive energy as possible, in the sense that $\sum_{n\leq x}r_{B,h}(n)^{2}\geq(\sum_{n\leq x}r_{B,h}(n))^{2}/x\asymp B(x)^{2h}/x$ by Cauchy–Schwarz. It will only be used in the proof of Proposition 2.2. In condition (iii), we roughly ask for the expectation of $r_{\mathscr{A},h}(n)$ for a random subset $\mathscr{A}\subseteq B$ of density $|\mathscr{A}\cap[1,x]|\approx f(x)$ to have a certain asymptotic behaviour, over a subset $\mathscr{S}\subseteq\mathbb{Z}_{\geq 0}$ that is usually an arithmetic progression. We observe that if $\omega<\beta$, then $f(x)/B(x)$ is necessarily asymptotic to a non-increasing function since $f(x)/B(x)\sim\sup_{y\geq x}f(y)/B(y)$ by [1, Theorem 1.5.3]. The only place we need to assume $f(x)/B(x)$ monotone is Proposition 2.2.
Theorem 1.4.
Let $B\subseteq\mathbb{Z}_{\geq 0}$ be a subset and $F$ be a regularly varying function satisfying conditions (i)(iii). Let $h\geq\max\{2H_{\textnormal{(ii)}}+1,H_{\textnormal{(iii)}}\}$ be an integer and suppose $F(n)\leq(1+o(1))r_{B,h}(n)$. If $F(n)/\log n\to\infty$ as $n\to\infty$, then there exists $A\subseteq B$ such that
$r_{A,h}(n)\sim\mathfrak{S}_{B,h}(n)F(n)\quad\text{for}\quad n\in\mathscr{S}.$
Moreover, if $F=\log$, then there exists $A\subseteq B$ such that
$r_{A,h}(n)\asymp\log n\quad\text{for}\quad n\in\mathscr{S}.$
In combination with Lemma 2.1, the proof of Theorem 1.4 also implies that for a given $F(x)=x^{\kappa}\psi(x)$, the set $A$ satisfies
$|A\cap[1,x]|\sim\frac{h\beta}{1+\kappa}C_{B,h,f}^{-1/h}(xF(x))^{1/h}.$ (1.4)
Notation.
Throughout the paper, we use the common asymptotic notation $\sim$, $o$, $O$, $\gg$, $\ll$, $\asymp$ with a subscript to indicate dependency on some parameter (omitting dependencies on $h$, $k$, $f$, $F$, and $B$). For a sequence of random variables $(X_{n})_{n\geq 1}$ in a probability space $\Omega$, and a real function $g$, we write $X_{n}\stackrel{{\scriptstyle\text{a.s.}}}{{\sim}}g(n)$ if $X_{n}(\omega)\sim g(n)$ for almost all $\omega\in\Omega$. Lastly, $e(\alpha):=e^{2\pi i\alpha}$.
2. Random subsets
Let $B\subseteq\mathbb{N}$ be a set satisfying conditions (i)(iii), and let $h\geq\max\{2H_{\textnormal{(ii)}}+1,H_{\textnormal{(iii)}}\}$ be an integer. We introduce a constant $c>0$, to be chosen later, and let
$f(x)=x^{\omega}\phi(x)$
be a regularly varying function such that $cf(x)\leq B(x)$ and $1/h\leq\omega\leq\beta$. We suppose that $f(x)/B(x)$ is non-increasing. [^2] Since $f(x)/B(x)$ is asymptotic to a non-increasing function by condition (iii), and we only care about asymptotics, we may substitute $f(x)$ by $B(x)(\sup_{y>x}f(y)/B(y))$ if needed. Now consider a probability space whose elements are subsets $0\in\mathscr{A}\subseteq B$, where
$\Pr(\mathbbm{1}_{\mathscr{A}}(n)=1)=\mathbb{E}(\mathbbm{1}_{\mathscr{A}}(n)):=c\frac{f(n)}{B(n)}\mathbbm{1}_{B}(n)\qquad(\forall n\in\mathbb{Z}_{\geq 1}),$ (2.1)
and where the variables $\mathbbm{1}_{\mathscr{A}}(n)$ are mutually independent $\{0,1\}$-random variables. This probability space has the feature that the counting function of $\mathscr{A}\subseteq B$ asymptotically matches $f$ with probability $1$:
Lemma 2.1.
$|\mathscr{A}\cap[1,x]|\stackrel{{\scriptstyle\textnormal{a.s.}}}{{\sim}}c\frac{\beta}{\omega}f(x)$.
Proof.
Since $B(x)$ is regularly varying, the function $b_{\lfloor x\rfloor}:=\inf\{y\in\mathbb{R}_{\geq 0}\mid B(y)\geq\lfloor x\rfloor\}$ is also regularly varying, with $b_{B(x)}\sim x$. Using the fact that compositions of regularly varying functions are also regularly varying, we can express $f(b_{\lfloor x\rfloor})=x^{\omega/\beta}\vartheta(x)$ for some slowly varying function $\vartheta(x)$. Notice that $B(x)^{\omega/\beta}\vartheta(B(x))=f(b_{B(x)})\sim f(x)$. By the strong law of large numbers, we have
$|\mathscr{A}\cap[1,x]|$ $\stackrel{{\scriptstyle\text{a.s.}}}{{\sim}}c\sum_{n\leq x}\frac{f(n)}{B(n)}\mathbbm{1}_{B}(n)=c\sum_{k\leq B(x)}\frac{f(b_{k})}{k}$
$\sim c\int_{1}^{B(x)}t^{\omega/\beta-1}\vartheta(t)\,\mathrm{d}t$
$\sim c\left(\int_{1/B(x)}^{1}u^{\omega/\beta-1}\frac{\vartheta(uB(x))}{\vartheta(B(x))}\,\mathrm{d}u\right)f(x).$
By Potter bounds (cf. BGT [1, Theorem 1.5.6(i)]), for any $\delta>0$ there exists $C=C_{\delta}>0$ such that, for large $x\geq x_{\delta}$ and $C/x\leq u\leq 1$, we have $\vartheta(ux)/\vartheta(x)\leq 2u^{-\delta}$. Selecting $\delta<\omega/2\beta$, we split the integral $\int_{1/B(x)}^{1}=\int_{1/B(x)}^{C/B(x)}+\int_{C/B(x)}^{1}$. Since by the definition of slow variation,
$u^{\omega/\beta-1}\frac{\vartheta(uB(x))}{\vartheta(B(x))}\mathbbm{1}_{(C/B(x),1]}\xrightarrow{x\to\infty}u^{\omega/\beta-1}\mathbbm{1}_{(0,1]},$
the dominated convergence theorem implies that
$\int_{C/B(x)}^{1}u^{\omega/\beta-1}\frac{\vartheta(uB(x))}{\vartheta(B(x))}\,\mathrm{d}u\to\int_{0}^{1}u^{\omega/\beta-1}\,\mathrm{d}u=\frac{\beta}{\omega}$
as $x\to\infty$. On the other hand, since $\vartheta(x)=x^{o(1)}$, the term $\int_{1/B(x)}^{C/B(x)}$ vanishes, completing the proof. ∎
2.1. Expectation of $r_{\mathscr{A},h}(n)$
We aim to compute the expectation of
$r_{\mathscr{A},h}(n)=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{Z}_{\geq 0}\\ x_{1}+\cdots+x_{h}=n\end{subarray}}\mathbbm{1}_{\mathscr{A}}(x_{1})\cdots\mathbbm{1}_{\mathscr{A}}(x_{h}).$
To facilitate this, we introduce auxiliary representation functions that lend themselves more readily to probabilistic analysis. Let $1\leq\ell\leq h$, and $c_{1},\ldots,c_{\ell}\in\mathbb{Z}_{\geq 1}$. A solution $(x_{1},\ldots,x_{h})$ to the equation $c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n$ is called exact if the $x_{i}$’s are pairwise distinct. The corresponding exact representation function (associated with $c_{1},\ldots,c_{\ell}$) is defined as
$\rho_{\mathscr{A},\ell}^{(c_{1},\ldots,c_{\ell})}(n):=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in\mathbb{Z}_{\geq 0}\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\\ x_{i}\text{ distinct}\end{subarray}}\mathbbm{1}_{\mathscr{A}}(x_{1})\cdots\mathbbm{1}_{\mathscr{A}}(x_{\ell}).$
For the special case where $c_{1}=\cdots=c_{\ell}=1$, we denote this function simply by $\rho_{\mathscr{A},\ell}(n)$.
Now, consider a solution $(x_{1},\ldots,x_{h})$ to $x_{1}+\cdots+x_{h}=n$ where $x_{1}=x_{2}$, but $x_{2},\ldots,x_{h}$ are pairwise distinct. In such a scenario, the tuple $(x_{2},\ldots,x_{h})$ forms an exact solution to the equation
$2x_{2}+x_{3}+\cdots+x_{h}=n.$
More generally, every non-exact solution of length $h$ corresponds to an exact solution of an equation of smaller length, where the coefficients sum to $h$. This leads to the decomposition:
$r_{\mathscr{A},h}(n)=\rho_{\mathscr{A},h}(n)+\sideset{}{{}^{\prime}}{\sum}_{\begin{subarray}{c}(c_{1},\ldots,c_{\ell})\end{subarray}}\rho_{\mathscr{A},\ell}^{(c_{1},\ldots,c_{\ell})}(n),$ (2.2)
where the summation runs over $11\leq\ell\leq h-$ 1 and all $(c_{1},\ldots,c_{\ell})\in\mathbb{Z}_{\geq 1}^{\ell}$ such that $c_{1}+\cdots+c_{\ell}=h$. The following proposition provides a key estimate for bounding contributions from non-exact solutions:
Proposition 2.2.
Let $h\geq 2H_{\textnormal{(ii)}}+1$, and $11\leq\ell\leq h-$ 1. Let $f(x)=x^{\omega}\phi(x)$ be a function of regular variation with $\omega\geq 1/h$. Let $c_{1}$, $\ldots$, $c_{\ell}\in\mathbb{Z}_{\geq 1}$ be such that $c_{1}+\cdots+c_{\ell}\leq C$. Then, there exists $\eta=\eta(h,f)>0$ such that
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in B\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots\frac{f(x_{\ell})}{B(x_{\ell})}\ll_{C}n^{-\eta}\frac{f(n)^{h}}{n},$
The proof of Proposition 2.2 is given in Subsection 2.2. The following two lemmas show that it suffices to consider exact solutions:
Lemma 2.3.
For $n\in\mathscr{S}$, we have
$\mathbb{E}(\rho_{\mathscr{A},h}(n))\sim c^{h}\,\mathfrak{S}_{B,h}(n)\,C_{B,h,f}\frac{f(n)^{h}}{n},$
and $\mathbb{E}(r_{\mathscr{A},h}(n))=(1+O(n^{-\eta}))\mathbb{E}(\rho_{\mathscr{A},h}(n))$, where $\eta$ is as in Proposition 2.2.
Proof.
By condition (iii) and Proposition 2.2, we have
$\mathbb{E}(\rho_{\mathscr{A},h}(n))=c^{h}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in B\\ x_{1}+\cdots+x_{h}=n\\ x_{i}\text{'s distinct}\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots\frac{f(x_{h})}{B(x_{h})}$
$=c^{h}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in B\\ x_{1}+\cdots+x_{h}=n\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots\frac{f(x_{h})}{B(x_{h})}-O\Bigg{(}c^{h}\,\sideset{}{{}^{\prime}}{\sum}_{\begin{subarray}{c}(c_{1},\ldots,c_{\ell})\end{subarray}}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in B\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots\frac{f(x_{\ell})}{B(x_{\ell})}\Bigg{)}$
$=(1+o(1))c^{h}\,\mathfrak{S}_{B,h}(n)\,C_{B,h,f}\frac{f(n)^{h}}{n},$
proving the first part. For the second part, by (2.2) we have
$\mathbb{E}(r_{\mathscr{A},h}(n))=\mathbb{E}(\rho_{\mathscr{A},h}(n))+\sideset{}{{}^{\prime}}{\sum}_{\begin{subarray}{c}(c_{1},\ldots,c_{\ell})\end{subarray}}\mathbb{E}(\rho^{(c_{1},\ldots,c_{\ell})}_{\mathscr{A},\ell}(n)),$
and for each $(c_{1},\ldots,c_{\ell})$ we have, by Proposition 2.2,
$\mathbb{E}(\rho^{(c_{1},\ldots,c_{\ell})}_{\mathscr{A},\ell}(n))$ $\ll\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in B\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots\frac{f(x_{\ell})}{B(x_{\ell})}\ll n^{-\eta}\frac{f(n)^{h}}{n},$
completing the proof. ∎
Lemma 2.4.
We have
$\mathbb{E}(r_{\mathscr{A},h-1}(n))\ll n^{-\eta}\frac{f(n)^{h}}{n}.$
Proof.
As in (2.2), we have
$\mathbb{E}(r_{\mathscr{A},h-1}(n))=\mathbb{E}(\rho_{\mathscr{A},h-1}(n))+\sideset{}{{}^{\prime\prime}}{\sum}_{\begin{subarray}{c}(c_{1},\ldots,c_{\ell})\end{subarray}}\mathbb{E}(\rho^{(c_{1},\ldots,c_{\ell})}_{\mathscr{A},\ell}(n)),$
where the sum runs over $21\leq\ell\leq h-$ 2 and the $(c_{1},\ldots,c_{\ell})\in\mathbb{Z}_{\geq 1}^{\ell}$ such that $c_{1}+\cdots+c_{\ell}=h-1$. As in the proof of Lemma 2.3, by Proposition 2.2 it follows that $\mathbb{E}(r_{\mathscr{A},h-1}(n))\ll n^{-\eta}f(n)^{h}/n$. ∎
2.2. Proof of Proposition 2.2
For $\alpha\in[0,1)$, $x\in\mathbb{R}_{\geq 0}$, let
$T(\alpha;x):=\sum_{n\leq x}\frac{f(n)}{B(n)}\mathbbm{1}_{B}(n)\,e(\alpha n),\qquad\widetilde{T}(\alpha;x):=\sum_{x/h^{2}\leq n\leq x}\frac{f(n)}{B(n)}\mathbbm{1}_{B}(n)\,e(\alpha n).$
Since every solution to $c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n$ has some $x_{i}\geq n/h^{2}$, we have
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in B\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots\frac{f(x_{\ell})}{B(x_{\ell})}\leq\sum_{\begin{subarray}{c}I\subseteq\{1,\ldots,\ell\}\\ I\neq\varnothing\end{subarray}}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in B\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\\ \forall i\in I,\,x_{i}\geq n/h^{2}\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots\frac{f(x_{\ell})}{B(x_{\ell})}$
By the orthogonality of the $e(n\alpha)$ and Hölder’s inequality, we have
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in B\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\\ \forall i\in I,\,x_{i}\geq n/h^{2}\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots$ $\frac{f(x_{\ell})}{B(x_{\ell})}=\int_{0}^{1}\Bigg{(}\prod_{i\in I}\widetilde{T}(c_{i}\alpha;n)\Bigg{)}\Bigg{(}\prod_{\begin{subarray}{c}j=1\\ j\notin I\end{subarray}}^{\ell}T(c_{j}\alpha;n)\Bigg{)}e(-n\alpha)\,\mathrm{d}\alpha$
$\leq\prod_{i\in I}\bigg{(}\int_{0}^{1}|\widetilde{T}(c_{i}\alpha;n)|^{\ell}\,\mathrm{d}\alpha\bigg{)}^{1/\ell}\prod_{\begin{subarray}{c}j=1\\ j\notin I\end{subarray}}^{\ell}\bigg{(}\int_{0}^{1}|T(c_{j}\alpha;n)|^{\ell}\,\mathrm{d}\alpha\bigg{)}^{1/\ell}$
$\ll_{C}\bigg{(}\int_{0}^{1}|\widetilde{T}(\alpha;n)|^{\ell}\,\mathrm{d}\alpha\bigg{)}^{|I|/\ell}\bigg{(}\int_{0}^{1}|T(\alpha;n)|^{\ell}\,\mathrm{d}\alpha\bigg{)}^{1-|I|/\ell},$
therefore
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in B\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\end{subarray}}\frac{f(x_{1})}{B(x_{1})}$ $\cdots\frac{f(x_{\ell})}{B(x_{\ell})}$ (2.3)
$\ll_{C}\sum_{j=1}^{\ell}\bigg{(}\int_{0}^{1}|\widetilde{T}(\alpha;n)|^{\ell}\,\mathrm{d}\alpha\bigg{)}^{j/\ell}\bigg{(}\int_{0}^{1}|T(\alpha;n)|^{\ell}\,\mathrm{d}\alpha\bigg{)}^{1-j/\ell}.$
Lemma 2.5.
Let $h\geq 2H_{\textnormal{(ii)}}$ be an integer. There exists $\delta=\delta(B,h,\omega)>0$ such that, for every $11\leq\ell\leq h-$ 1,
$\int_{0}^{1}|T(\alpha;n)|^{\ell}\,\mathrm{d}\alpha\ll n^{\max\{h\omega-1-\delta,0\}+o(1)}$
and
$\int_{0}^{1}|\widetilde{T}(\alpha;n)|^{\ell}\,\mathrm{d}\alpha\ll n^{\max\{h\omega-1,0\}-\delta+o(1)}.$
Proof.
Let $g(\alpha;x):=\sum_{n\leq x}\mathbbm{1}_{B}(n)e(n\alpha)$. Since $f(2x)/B(2x)\asymp f(x)/B(x)$ by condition (i), by partial summation we have
$T(\alpha$ $;x)=\sum_{j\leq\frac{\log x}{\log 2}}\sum_{x/2^{j+1}<n\leq x/2^{j}}\frac{f(n)}{B(n)}\mathbbm{1}_{B}(n)e(n\alpha)$
$\ll\sum_{j\leq\frac{\log x}{\log 2}}\bigg{(}\sup_{x/2^{j+1}<y\leq x/2^{j}}|g(\alpha;y)|\bigg{)}\Bigg{(}\frac{f(2^{-j}x)}{B(2^{-j}x)}+\sum_{x/2^{j+1}<n\leq x/2^{j}}\bigg{|}\frac{f(n)}{B(n)}-\frac{f(n+1)}{B(n+1)}\bigg{|}\Bigg{)}.$
The function $f(x)/B(x)$ is non-increasing, so we have $\sum_{x/2^{j+1}<n\leq x/2^{j}}\big{|}\frac{f(n)}{B(n)}-\frac{f(n+1)}{B(n+1)}\big{|}\ll f(2^{-j}x)/B(2^{-j}x)$. It follows that
$T(\alpha;x)$ $\ll(\log x)\max_{j\leq\frac{\log x}{\log 2}}\bigg{(}\bigg{(}\sup_{x/2^{j+1}<y\leq x/2^{j}}|g(\alpha;y)|\bigg{)}\frac{f(2^{-j}x)}{B(2^{-j}x)}\bigg{)}.$
So, for $\ell^{*}:=\max\{\ell,2H_{\textnormal{(ii)}}\}$, we have
$\int_{0}^{1}|T(\alpha;n)|^{\ell}\,\mathrm{d}\alpha$ $\leq\bigg{(}\int_{0}^{1}|T(\alpha;n)|^{\ell^{*}}\,\mathrm{d}\alpha\bigg{)}^{\ell/\ell^{*}}$
$\ll(\log n)^{\ell}\max_{j\leq\frac{\log n}{\log 2}}\bigg{(}\bigg{(}\sup_{n/2^{j+1}<y\leq n/2^{j}}\int_{0}^{1}|g(\alpha;y)|^{\ell^{*}}\,\mathrm{d}\alpha\bigg{)}^{\ell/\ell^{*}}\frac{f(2^{-j}n)^{\ell}}{B(2^{-j}n)^{\ell}}\bigg{)}.$
By Parseval’s identity and condition (ii), for $n/2^{j+1}<y\leq n/2^{j}$ we have
$\int_{0}^{1}|g(\alpha;y)|^{\ell^{*}}\,\mathrm{d}\alpha$ $=\int_{0}^{1}\bigg{|}\sum_{m\leq y}\mathbbm{1}_{B}(m)e(m\alpha)\bigg{|}^{\ell^{*}}\,\mathrm{d}\alpha$
$\leq B(y)^{\ell^{*}-2H_{\textnormal{(ii)}}}\int_{0}^{1}\bigg{|}\sum_{m\leq y}\mathbbm{1}_{B}(m)e(m\alpha)\bigg{|}^{2H_{\textnormal{(ii)}}}\,\mathrm{d}\alpha$
$\leq B(y)^{\ell^{*}-2H_{\textnormal{(ii)}}}\sum_{m\leq H_{\textnormal{(ii)}}y}r_{B,H_{\textnormal{(ii)}}}(m)^{2}$
$\ll\frac{B(y)^{\ell^{*}+o(1)}}{y}\asymp\frac{B(2^{-j}n)^{\ell^{*}+o(1)}}{2^{-j}n},$
and thus, as $B(x)^{o(1)}=x^{o(1)}$ and $f(x)=x^{\omega+o(1)}$,
$\int_{0}^{1}|T(\alpha;n)|^{\ell}\,\mathrm{d}\alpha\ll(\log n)^{\ell}\max_{j\leq\frac{\log n}{\log 2}}\bigg{(}\frac{f(2^{-j}n)^{\ell+o(1)}}{(2^{-j}n)^{\ell/\ell^{*}}}\bigg{)}=n^{\max\{\ell\omega-\frac{\ell}{\ell^{*}},0\}+o(1)}.$ (2.4)
By a similar argument to the one above used to bound $T$, we obtain
$\int_{0}^{1}|\widetilde{T}(\alpha;n)|^{\ell}\,\mathrm{d}\alpha$ $\leq\bigg{(}\int_{0}^{1}|\widetilde{T}(\alpha;n)|^{\ell}\,\mathrm{d}\alpha\bigg{)}^{\ell/\ell^{*}}$
$\ll\bigg{(}\sup_{n/h^{2}\leq y\leq n}\int_{0}^{1}|g(\alpha;y)|^{\ell}\,\mathrm{d}\alpha\bigg{)}^{\ell/\ell^{*}}\frac{f(n)^{\ell}}{B(n)^{\ell}}$
$\ll\bigg{(}\frac{B(n)^{\ell^{*}+o(1)}}{n}\bigg{)}^{\ell/\ell^{*}}\frac{f(n)^{\ell}}{B(n)^{\ell}}$
$=\frac{f(n)^{\ell+o(1)}}{n^{\ell/\ell^{*}}}=n^{\ell\omega-\frac{\ell}{\ell^{*}}+o(1)}.$ (2.5)
Using (2.4), (2.5), to finish the proof, it suffices to show that for every $11\leq\ell\leq h-$ 1, if $\ell\omega-\frac{\ell}{\ell^{*}}\geq 0$ then $\ell\omega-\frac{\ell}{\ell^{*}}<h\omega-1$. This is trivial for $\ell\geq 2H_{\textnormal{(ii)}}$ (since $\ell^{*}=\ell$), so suppose $\ell<2H_{\textnormal{(ii)}}$. We have $\ell\omega-\frac{\ell}{\ell^{*}}=\ell(\omega-\frac{1}{2H_{\textnormal{(ii)}}})$, so we must have $\omega\geq\frac{1}{2H_{\textnormal{(ii)}}}$; thus,
$\ell\bigg{(}\omega-\frac{1}{2H_{\textnormal{(ii)}}}\bigg{)}\leq(2H_{\textnormal{(ii)}}-1)\bigg{(}\omega-\frac{1}{2H_{\textnormal{(ii)}}}\bigg{)}$ $=2H_{\textnormal{(ii)}}\omega-\bigg{(}\omega-\frac{1}{2H_{\textnormal{(ii)}}}\bigg{)}-1$
$<h\omega-1,$
the last line following since $h\geq 2H_{\textnormal{(ii)}}+1$. ∎
Applying Lemma 2.5 to (2.3), we obtain
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in B\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots\frac{f(x_{\ell})}{B(x_{\ell})}$ $\ll_{C}\big{(}n^{\max\{h\omega-1,0\}-\delta+o(1)}\big{)}^{1/\ell}\big{(}n^{\max\{h\omega-1-\delta,0\}+o(1)}\big{)}^{1-1/\ell}$
$\ll n^{\max\{h\omega-1,0\}-\delta/\ell+o(1)}.$
Since $\omega\geq 1/h$ and $f(n)^{h}/n=x^{h\omega-1+o(1)}$, we have
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in B\\ c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n\end{subarray}}\frac{f(x_{1})}{B(x_{1})}\cdots\frac{f(x_{\ell})}{B(x_{\ell})}\ll_{C}n^{-\delta/\ell+o(1)}\frac{f(n)^{h}}{n},$
concluding the proof of Proposition 2.2. $\square$
3. Proof of Theorem 1.4
For the remainder of this section, fix $h\geq\max\{2H_{\textnormal{(ii)}}+1,H_{\textnormal{(iii)}}\}$, and assume
$n\in\mathscr{S},\quad n\to\infty.$
Consider the function $F(x)=x^{\kappa}\psi(x)$, where $\psi(x)$ is a slowly varying function, and $\kappa$ is a real number satisfying $10\leq\kappa\leq h-$ 1. Define
$f(x):=(xF(x))^{1/h}=x^{(1+\kappa)/h}\psi(x)^{1/h}.$
It is sufficient to work with $f(x)$ under the conditions
$\frac{f(x)}{(x\log x)^{1/h}}\xrightarrow{x\to\infty}\infty,\quad\text{and}\quad cf(x)\leq x.$
Using Lemma 2.3, the proof of Theorem 1.4 reduces to showing that, in the probability space defined in Section 2, we have
$r_{\mathscr{A},h}(n)\stackrel{{\scriptstyle\textnormal{a.s.}}}{{\sim}}\mathbb{E}(r_{\mathscr{A},h}(n)).$
Since an event of probability one is non-empty, this implies the existence of a subset $A\subseteq B$ such that $r_{A,h}(n)\sim\mathbb{E}(r_{\mathscr{A},h}(n))$. To proceed, we begin with two essential concentration lemmas.
3.1. Concentration of boolean polynomials
Let $n\in\mathbb{Z}_{\geq 1}$, and consider independent (not necessarily identically distributed) random variables $v_{1},\ldots,v_{n}$, each taking values in $\{0,1\}$. A boolean polynomial is a multivariate polynomial expressed as
$Y(v_{1},\ldots,v_{n})=\sum_{i}c_{i}I_{i}\in\mathbb{R}[v_{1},\ldots,v_{n}],$
where each $I_{i}$ is a monomial in the variables $v_{k}$. We classify such polynomials as follows:
  • Positive: If $c_{i}>0$ for all $i$,
  • Simple: If the largest exponent of any $v_{k}$ in any monomial is at most $1$,
  • Homogeneous: If all monomials have the same total degree,
  • Normal: If $0\leq c_{i}\leq 1$ for all $i$, and the constant term of $Y$ is zero.
For a non-empty multiset [^3] A multiset is a set allowing repeated elements. $S\subseteq\{v_{1},\ldots,v_{n}\}$, define
$\partial_{S}:=\prod_{v\in S}\partial_{v},$
where $\partial_{v}$ denotes the partial derivative with respect to $v$. For example, if $S=\{1,1,2\}$, then
$\partial_{S}(v_{1}^{3}v_{2}v_{3}+3v_{1}^{5})=6v_{1}v_{3}.$
We further define the quantities:
$\mathbb{E}_{j}(Y):=\max_{\begin{subarray}{c}S\subseteq\{v_{1},\ldots,v_{n}\}\\ \text{multiset},|S|=j\end{subarray}}\mathbb{E}(\partial_{S}Y),\quad\text{and}\quad\mathbb{E}^{\prime}(Y):=\max_{j\geq 1}\mathbb{E}_{j}(Y).$
Two key concentration results will be applied:
Theorem 3.1 (Kim–Vu [9]).
Let $d\geq 1$, and suppose $Y(v_{1},\ldots,v_{n})$ is a positive, simple boolean polynomial of degree $d$. Define $E^{\prime}:=\mathbb{E}^{\prime}(Y)$ and $E:=\max\{\mathbb{E}(Y),E^{\prime}\}$. Then, for any $\lambda\geq 1$,
$\Pr\big{(}|Y-\mathbb{E}(Y)|>8^{d}\sqrt{d!}\,\lambda^{d-1/2}(E^{\prime}E)^{1/2}\big{)}\ll_{d}n^{d-1}e^{-\lambda}.$
In applications, we set $\lambda=(d+1)\log n$. This result is most useful when $1\ll\mathbb{E}^{\prime}\ll\mathbb{E}/(\log n)^{2d}$. The next lemma deals with the case where $\mathbb{E}$ is small.
Theorem 3.2 (Vu [17, Theorem 1.4]).
Let $d\geq 2$, and let $Y(v_{1},\ldots,v_{n})=\sum_{i}c_{i}I_{i}$ be a simple, homogeneous, normal boolean polynomial of degree $d$. For any $\alpha,\gamma>0$, there exists a constant $K=K(d,\alpha,\gamma)$ such that if $\mathbb{E}_{1}(Y),\ldots,\mathbb{E}_{d-1}(Y)\leq n^{-\alpha}$, then for any $0<\lambda\leq\mathbb{E}(Y)$,
$\Pr(|Y-\mathbb{E}(Y)|\geq(\lambda\mathbb{E}(Y))^{1/2})\leq 2he^{-\lambda/16dK}+n^{-\gamma}.$
Remark.
The random variable $r_{\mathscr{A},h}(n)$ can be viewed as a boolean polynomial:
$r_{\mathscr{A},h}(n)=Y(v_{1},\ldots,v_{n}),$
where $v_{i}$ are independent $\{0,1\}$-random variables with $\Pr(v_{i}=1)=\mathbb{E}(\mathbbm{1}_{\mathscr{A}}(i))$. The expected values of its partial derivatives are bounded above by
$\mathbb{E}(r_{\mathscr{A},h-\ell}(n-k)),\quad 1\leq k\leq n,$
for $1\leq\ell\leq h$.
3.2. Theorem 1.4: Case $\kappa>0$
Set $c:=C_{B,h,f}^{-1/h}$. We will now apply Kim–Vu’s inequality to the function $\rho_{\mathscr{A},h}(n)$. Using the notation of Theorem 3.1, we have
$E$ $=\mathbb{E}(\rho_{\mathscr{A},h}(n)),\quad\text{ and }\quad E^{\prime}\leq 1+\max\limits_{1\leq\ell\leq h-1}\max\limits_{k\leq n}\mathbb{E}(r_{\mathscr{A},\ell}(k))\leq 1+\max\limits_{k\leq n}\mathbb{E}(r_{\mathscr{A},h-1}(k)).$
By Lemmas 2.3 and 2.4, and setting $\lambda=(h+1)\log n$, we find
$\lambda^{h-\frac{1}{2}}(E^{\prime}E)^{1/2}$ $\ll(h+1)^{h}(\log n)^{h}\bigg{(}\frac{E^{\prime}}{E}\bigg{)}^{1/2}\,E$
$\ll(\log n)^{h}\Bigg{(}\frac{n^{\max\{0,\kappa-\eta\}+o(1)}}{n^{\kappa+o(1)}}\Bigg{)}^{1/2}E=\frac{(\log n)^{h}}{n^{\min\{\kappa,\eta\}/2+o(1)}}\,E.$
This leads to the bound:
$\Pr\Bigg{(}\big{|}\rho_{\mathscr{A},h}(n)-\mathbb{E}(\rho_{\mathscr{A},h}(n))\big{|}\geq\frac{8^{h}\sqrt{h!}}{n^{\min\{\kappa,\eta\}/2+o(1)}}\,\mathbb{E}(\rho_{\mathscr{A},h}(n))\Bigg{)}\ll n^{-2}.$ (3.1)
Thus, the Borel–Cantelli lemma implies that $\rho_{\mathscr{A},h}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\sim}}\mathbb{E}(\rho_{\mathscr{A},h}(n))\sim\mathfrak{S}_{B,h}(n)F(n)$.
To handle the contribution of non-exact solutions, observe that by (2.2), it suffices to show
$\rho_{\mathscr{A},\ell}^{(c_{1},\ldots,c_{\ell})}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{=}}o(F(n)),$
for every $11\leq\ell\leq h-$ 1 and for equations $c_{1}x_{1}+\cdots+c_{\ell}x_{\ell}=n$ with $c_{1}+\cdots+c_{\ell}\leq h$. Applying Kim–Vu’s inequality to the boolean polynomial $\rho_{\mathscr{A},\ell}^{(c_{1},\ldots,c_{\ell})}(n)$, Lemma 2.3 gives
$E,\,E^{\prime}$ $\ll 1+n^{-\eta}\frac{f(n)^{h}}{n}=n^{\max\{0,\kappa-\eta\}+o(1)}.$
From this, we deduce
$\Pr\left(|\rho_{\mathscr{A},\ell}^{(c_{1},\ldots,c_{\ell})}(n)-\mathbb{E}(\rho_{\mathscr{A},\ell}^{(c_{1},\ldots,c_{\ell})}(n))|\geq 8^{h}\sqrt{h!}\,(\log n)^{h}\,n^{\max\{0,\kappa-\eta\}+o(1)}\right)\ll n^{-2}.$
Thus, by the Borel–Cantelli lemma,
$\rho_{\mathscr{A},\ell}^{(c_{1},\ldots,c_{\ell})}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\ll}}\mathbb{E}(\rho_{\mathscr{A},\ell}^{(c_{1},\ldots,c_{\ell})}(n))+n^{\max\{0,\kappa-\eta\}+o(1)}=o(F(n)).$
This completes the proof. $\square$
3.3. $\delta$-small and $\delta$-normal solutions
Following Vu [16], to estimate $r_{\mathscr{A},h}(n)$, we divide the solutions being counted into two types, based on a parameter $\delta$, as follows: for $10<\delta<$ 1, define
$r^{(\delta\textnormal{-small})}_{\mathscr{A},h}(n)$ $:=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{Z}_{\geq 0}\\ x_{1}+\cdots+x_{h}=n\\ \exists j\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{j}<n^{\delta}\end{subarray}}\mathbbm{1}_{\mathscr{A}}(x_{1})\cdots\mathbbm{1}_{\mathscr{A}}(x_{h}),$ (3.2)
$r^{(\delta\textnormal{-normal})}_{\mathscr{A},h}(n)$ $:=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{Z}_{\geq 0}\\ x_{1}+\cdots+x_{h}=n\\ x_{1},\ldots,x_{h}\geq n^{\delta}\end{subarray}}\mathbbm{1}_{\mathscr{A}}(x_{1})\cdots\mathbbm{1}_{\mathscr{A}}(x_{h}),$
so that $r_{\mathscr{A},h}(n)=r^{(\delta\textnormal{-small})}_{\mathscr{A},h}(n)+r^{(\delta\textnormal{-normal})}_{\mathscr{A},h}(n)$. The analogous quantities $\rho^{(\delta\textnormal{-small})}_{\mathscr{A},h}(n)$, $\rho^{(\delta\textnormal{-normal})}_{\mathscr{A},h}(n)$ are defined similarly. We aim to show that $\delta$-small solution contribute negligibly, on average.
Lemma 3.3.
Let $\eta$ be as in Proposition 2.2. For every $0<\delta<\eta$, we have
$\mathbb{E}\big{(}r^{(\delta\textnormal{-small})}_{\mathscr{A},h}(n)\big{)}\ll c^{h}\,n^{\beta\delta-\eta+o(1)}\,\frac{f(n)^{h}}{n}.$
Proof.
We begin by bounding the expectation:
$\mathbb{E}(r^{(\delta\textnormal{-small})}_{\mathscr{A},h}(n))$ $\leq\sum_{k\leq n^{\delta}}\mathbb{E}\big{(}r_{\mathscr{A},h-1}(n-k)\mathbbm{1}_{\mathscr{A}}(k)\big{)}$
[…]
For $k\leq n^{\delta}$, we apply Lemma 2.4 to bound the conditional expectation:
[…]
$\ll c^{h-1}\,n^{-\eta}\,\frac{f(n)^{h}}{n}.$
Substituting this back, we find
[…]
Since $f(k)$ and $B(k)$ are regularly varying, the sum is asymptotically proportional to $f(n^{\delta})$ (by the same argument used in Lemma 2.1). Therefore, we obtain
$\mathbb{E}(r^{(\delta\textnormal{-small})}_{\mathscr{A},h}(n))\asymp c^{h}\,n^{-\eta}f(n^{\delta})\,\frac{f(n)^{h}}{n}=c^{h}\,n^{\beta\delta-\eta+o(1)}\,\frac{f(n)^{h}}{n}.\qed$
3.4. Disjoint families of representations
Let $\widehat{r}_{\mathscr{A},\ell}(n)$ denote the maximum size of a disjoint family of representations $R=(x_{1},\ldots,x_{\ell})\in\mathscr{A}^{\ell}$ to $x_{1}+\cdots+x_{\ell}=n$. In other words, $\widehat{r}_{\mathscr{A},\ell}(n)=|\mathcal{M}|$, where $\mathcal{M}$ is a maximal disjoint family (abbreviated maxdisfam) of such representations. By definition, for every representation $R$, there exists a representation $S\in\mathcal{M}$ such that $S\cap R\neq\varnothing$. This implies the inequality:
$r_{\mathscr{A},\ell}(n)\leq\sum_{\begin{subarray}{c}k\in S\\ S\in\mathcal{M}\end{subarray}}r_{\mathscr{A},\ell-1}(n-k)\leq\ell!\,\widehat{r}_{\mathscr{A},\ell}(n)\cdot\left(\max_{k\leq n}r_{\mathscr{A},\ell-1}(k)\right).$ (3.3)
To indicate that a representation function refers to the size of a maximum disjoint family of representations, we add a “ $\widehat{\ }$ ” to its notation. For example, $\widehat{\rho}^{(\delta\text{-small})}_{\mathscr{A},\ell}$ denotes the size of the largest maxdisfam of $\delta$-small exact representations.
Lemma 3.4.
For every $2\leq\ell\leq h$, we have
$\mathrm{R}_{\mathscr{A},\ell}(n)\ll\widehat{\mathrm{R}}_{\mathscr{A},\ell}(n)\left(\max_{k\leq n}\widehat{r}_{\mathscr{A},\ell-1}(k)\right)\cdots\left(\max_{k\leq n}\widehat{r}_{\mathscr{A},2}(k)\right),$
where $\mathrm{R}=r$, $\rho$, $\rho^{(\delta\textnormal{-small})}$.
Proof.
Starting with (3.3) applied to $\mathrm{R}$, the same bound applies recursively to $r_{\mathscr{A},\ell-t}$ for $21\leq t\leq\ell-$ 2. This gives:
$\mathrm{R}_{\mathscr{A},\ell}(n)\ll\widehat{\mathrm{R}}_{\mathscr{A},\ell}(n)\left(\max_{k\leq n}\widehat{r}_{\mathscr{A},\ell-1}(k)\right)\cdots\left(\max_{k\leq n}\widehat{r}_{\mathscr{A},3}(k)\right)\left(\max_{k\leq n}r_{\mathscr{A},2}(k)\right).$
Since $r_{\mathscr{A},2}(k)\leq 2\widehat{r}_{\mathscr{A},2}(k)$, the conclusion follows. ∎
Lemma 3.4 will be used in conjunction with the following lemma [5, Lemma 1]:
Lemma 3.5 (Disjointness lemma).
Let =$\mathscr{E}=\{E_{1},E_{2},\ldots\}$ be a family of events, and define $S:=\sum_{E\in\mathscr{E}}\mathbbm{1}_{E}$. If $\mathbb{E}(S)<\infty$, then for every $k\in\mathbb{Z}_{\geq 1}$, we have:
[…]
Proof.
$\sum_{\begin{subarray}{c}\mathcal{J}\subseteq\mathscr{E}\textnormal{ disfam}\\ |\mathcal{J}|=k\end{subarray}}\Pr\bigg{(}\bigwedge_{E\in\mathcal{J}}E\bigg{)}=\sum_{\begin{subarray}{c}\mathcal{J}\subseteq\mathscr{E}\textnormal{ disfam}\\ |\mathcal{J}|=k\end{subarray}}\prod_{E\in\mathcal{J}}\Pr(E)$ $\leq\frac{1}{k!}\bigg{(}\sum_{E\in\mathscr{E}}\Pr(E)\bigg{)}^{k}=\frac{\mathbb{E}(S)^{k}}{k!}.\qed$
Remark.
If a maxdisfam of size greater than $k$ exists, then a disjoint family of size $k$ must also exist. Therefore, we apply Lemma 3.5 in the following form: since $k!\geq k^{k}e^{-k}$,
Pr(maxdisfam of $\Pr(\exists\text{maxdisfam of size}\geq x)$ $\leq\bigg{(}\frac{e\,\mathbb{E}}{\lceil x\rceil}\bigg{)}^{\lceil x\rceil}\qquad\text{(for real $x\geq 1$)}$
$\leq\bigg{(}\frac{e\,\mathbb{E}}{x}\bigg{)}^{x}$(for real $x\geq 1+\mathbb{E}$).
3.5. Theorem 1.4: Case $\kappa=0$
Define $c:=C_{B,h,f}^{-1/h}$. In this case, we have $\frac{f(x)^{h}}{x}=F(x)=x^{o(1)}$. Choose $0<\delta<\eta$, where $\eta$ is as given in Proposition 2.2. We will apply Theorem 3.2 to $\rho_{\mathscr{A},h}^{(\delta\text{-normal})}(n)$. After that, By (2.2) and Lemma 3.3, it will remain to show that $\rho^{(\delta\text{-small})}_{\mathscr{A},h}(n)$ and $r_{\mathscr{A},h-1}(n)$ are almost surely $O(1)$.
The function $\rho_{\mathscr{A},h}^{(\delta\text{-normal})}(n)$ is a homogeneous, simple boolean polynomial of degree $h$. Using the notation of Theorem 3.2, the partial derivatives of $\rho_{\mathscr{A},h}^{(\delta\text{-normal})}(n)$ are, for $11\leq j\leq h-$ 1, bounded by
$\mathbb{E}_{j}(\rho_{\mathscr{A},h}^{(\delta\text{-normal})}(n))\leq\max_{n^{\delta}\leq k\leq n}\mathbb{E}(r_{\mathscr{A},h-j}(k))$ $\ll_{c}\max_{n^{\delta}\leq k\leq n}k^{-\eta}\frac{f(k)^{h}}{k}\ll n^{-\eta\delta+o(1)},$
where the second step uses Lemma 2.4. Since each monomial of $\rho_{\mathscr{A},h}(n)$ appears at most $h!$ times, the function $\frac{1}{h!}\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)$ is a normal boolean polynomial. Taking $0<\alpha<\eta\delta$, $\gamma=2$, and $K=K(h,\alpha,\gamma)$ in Theorem 3.2, we get:
$\Pr\left(|\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)-\mathbb{E}(\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n))|\geq(h!\lambda\,\mathbb{E}(\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)))^{1/2}\right)\ll n^{-2},$
where we set $\lambda:=32hK\log n$. If $F(n)/\log n\to\infty$ as $x\to\infty$, then Lemma 3.3 implies $(\lambda\mathbb{E}(\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)))^{1/2}=o(\mathbb{E}(\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)))$. Thus, by the Borel–Cantelli lemma, we have $\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\sim}}\mathbb{E}(\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n))$. For $F=c\log$ with $c>1$ chosen sufficiently large to ensure $\mathbb{E}(\rho_{\mathscr{A},h}(n))>32hK\log n$, it follows that $\rho^{(\delta\text{-normal})}_{\mathscr{A},h}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\asymp}}\log n$.
To bound $\rho^{(\delta\text{-small})}_{\mathscr{A},h}(n)$, we apply Lemma 3.4. By Lemma 2.4, we have:
$\mathbb{E}(r_{\mathscr{A},\ell}(n))\ll n^{-\eta+o(1)},\qquad 1\leq\ell\leq h-1,$
and by Lemma 3.3 we have $\mathbb{E}(\rho^{(\delta\text{-small})}_{\mathscr{A},h}(n))\ll n^{-\eta+\beta\delta+o(1)}$. By the disjointness lemma 3.5,
$\Pr(\widehat{r}_{\mathscr{A},\ell}(n)\geq T)\leq\bigg{(}\frac{e}{T}\bigg{)}^{T}n^{-T\eta+o(1)}\ll n^{-2}$
for large $T>0$. A similar bound holds for $\widehat{\rho}^{(\delta\text{-small})}_{\mathscr{A},h}(n)$, since $-\eta+\beta\delta>0$. Thus, by the Borel–Cantelli lemma, we have $\widehat{r}_{\mathscr{A},\ell}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\ll}}1$ for $11\leq\ell\leq h-$ 1 and $\widehat{\rho}^{(\delta\text{-small})}_{\mathscr{A},h}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\ll}}1$. For each $12\leq\ell\leq h-$ 1, it follows that $\max_{k\leq n}\widehat{r}_{\mathscr{A},\ell}(k)\stackrel{{\scriptstyle\text{a.s.}}}{{\ll}}1$. Substituting this into Lemma 3.4, we conclude:
$\rho^{(\delta\text{-small})}_{\mathscr{A},h}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\ll}}\widehat{\rho}^{(\delta\text{-small})}_{\mathscr{A},h}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\ll}}1.$
Finally, to bound $r_{\mathscr{A},h-1}(n)$, if $h=2$ then $r_{\mathscr{A},h-1}(n)\leq 1$ trivially. For $h\geq 3$, we apply Lemma 3.4 again to obtain $r_{\mathscr{A},h-1}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\ll}}\widehat{r}_{\mathscr{A},h-1}(n)\stackrel{{\scriptstyle\text{a.s.}}}{{\ll}}1$. This completes the proof. $\square$
4. Waring subbases
In this section, we prove Theorem 1.1 by showing that $\mathbb{N}^{k}$ satisfies conditions (i)(iii). Clearly $\mathbb{N}^{k}$ satisfies condition (i), with $\beta=1/k$. As noted by Wooley [20, Equation (1.2)], $\mathbb{N}^{k}$ satisfies condition (ii) for $H_{\textnormal{(ii)}}\leq\frac{1}{2}k^{2}(\log k+\log\log k+O(1))$. More precisely,
$\int_{0}^{1}\bigg{|}\sum_{1\leq n\leq x}e(\alpha n^{k})\bigg{|}^{2\ell}\,\mathrm{d}\alpha\ll x^{2\ell-k+o(1)}$ (4.1)
for every $\ell\geq\frac{1}{2}k^{2}(\log k+\log\log k+O(1))$. By Parseval’s identity, we have
$\sum_{n\leq x}r_{\mathbb{N}^{k},\ell}(n)^{2}$ $\leq\int_{0}^{1}\bigg{|}\sum_{n\leq x^{1/\ell}}e(\alpha n^{k})\bigg{|}^{2h}\,\mathrm{d}\alpha\ll x^{2\ell/k-1+o(1)},$
which is the shape of condition (ii).
4.1. Waring satisfies condition (iii)
The first result we need, due to Wooley, is equivalent to condition (iii) restricted to functions of the form $x^{\omega}$; that is, an estimate for the number of solutions to $x_{1}^{k}+\cdots+x_{h}^{k}=n$ counted with weight $(x_{1}\cdots x_{h})^{\omega}$ for some $\omega>0$. By Theorem 1.4, this directy implies Theorem 1.1 (i).
Lemma 4.1 (Wooley [20, Theorem 1.1]).
There exists an integer $H_{W}\leq k^{2}(\log k+\log\log k+O(1))$ such that the following holds: Let $h\geq H_{W}$. For any $\omega\geq 1/h$,
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{N}^{k}\\ x_{1}+\cdots+x_{h}=n\end{subarray}}(x_{1}\cdots x_{h})^{\omega-\frac{1}{k}}=\mathfrak{S}_{k,h}(n)\frac{1}{k^{h}}\frac{\Gamma(\omega)^{h}}{\Gamma(h\omega)}\,n^{\omega h-1}+O(n^{\omega h-1-\tau})$
for some $\tau=\tau(k,h)>0$.
To prove Theorem 1.1 (ii), we need a second result, due to Vu, on an upper bound to the number of solutions $x_{1}^{k}+\cdots+x_{h}^{k}=n$ for $x_{j}$’s restricted to some box.
Lemma 4.2 (Vu [16, Lemma 2.1]).
There exists an integer $H_{V}\ll 8^{k}k^{2}$ such that the following holds: Let $h\geq H_{V}$. Given $P_{1}$, $\ldots$, $P_{h}\geq 1$, the number of solutions to
$x_{1}+\cdots+x_{h}=n$
with $x_{j}\in\mathbb{N}^{k}\cap[1,P_{j}]$ ($1\leq j\leq h$) is
$O\bigg{(}\frac{1}{n}(P_{1}\cdots P_{h})^{\frac{1}{k}}+(P_{1}\cdots P_{h})^{\frac{1}{k}-\frac{1}{h}-\delta}\bigg{)},$
for some $\delta=\delta(k,h)>0$.
Let $\phi$ be a slowly varying function. By uniform convergence (cf. BGT [1, Theorem 1.2.1]), given $10<\mu<$ 1, for every $\varepsilon>0$ there exists $x_{\mu,\varepsilon}\in\mathbb{R}$ such that, for every $x\geq x_{\mu,\varepsilon}$,
$\bigg{|}\frac{\phi(\lambda x)}{\phi(x)}-1\bigg{|}<\varepsilon,\qquad\forall\lambda\in[\mu,1].$
Taking $\mu=1/j$, $j\in\mathbb{Z}_{\geq 1}$, for large $x$ we define $\epsilon(x):=j$ for $x\in[x_{\frac{1}{j},\frac{1}{j}},x_{\frac{1}{j+1},\frac{1}{j+1}})$. This defines a non-decreasing function $\epsilon(x)\to\infty$ such that
$\frac{\phi(y)}{\phi(x)}\to 1\text{ uniformly for }y\in\bigg{[}\frac{x}{\epsilon(x)},x\bigg{)}.$ (4.2)
Lemma 4.3.
Let $h\geq\max\{H_{V},H_{W}\}$. We have:
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{N}^{k}\\ x_{1}+\cdots+x_{h}=n\end{subarray}}\frac{f(x_{1})}{x_{1}^{1/k}}\cdots\frac{f(x_{h})}{x_{h}^{1/k}}\sim\frac{1}{k^{h}}\frac{\Gamma(\omega)^{h}}{\Gamma(h\omega)}\mathfrak{S}_{k,h}(n)\frac{f(n)^{h}}{n}.$
Proof.
Let $\epsilon(x)$ be as in (4.2). Start by separating the sum into
$S_{1}:=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{N}^{k}\\ x_{1}+\cdots+x_{h}=n\\ \exists j\,|\,x_{j}<n/\epsilon(n)\end{subarray}}\frac{f(x_{1})}{x_{1}^{1/k}}\cdots\frac{f(x_{h})}{x_{h}^{1/k}},\qquad S_{2}:=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{N}^{k}\\ x_{1}+\cdots+x_{h}=n\\ \forall j,\,x_{j}\geq n/\epsilon(n)\end{subarray}}\frac{f(x_{1})}{x_{1}^{1/k}}\cdots\frac{f(x_{h})}{x_{h}^{1/k}}.$
We start with $S_{1}$. Let $\mathcal{P}$ be the set of all $h$-tuples $\mathbf{p}=(P_{1},\ldots,P_{h})$ with $P_{1}\in\{1,2,4\ldots,2^{L}\}$, and $P_{j}\in\{1,2,4,\ldots,2^{J}\}$ ($2\leq j\leq h$), where $L$ (resp. $J$) is the smallest integer for which $2^{L}\geq n/\varepsilon(n)$ (resp. $2^{J}\geq n$), and write
$\sigma_{\mathbf{p}}:=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{N}^{k}\\ x_{1}+\cdots+x_{h}=n\\ \frac{P_{j}}{2}-\frac{1}{2}\leq x_{j}<P_{j},\ \forall j\end{subarray}}\frac{f(x_{1})}{x_{1}^{1/k}}\cdots\frac{f(x_{h})}{x_{h}^{1/k}}$
(where “$f(0)/0=1$”). We have $S_{1}\leq\sum_{\mathbf{p}\in\mathcal{P}}\sigma_{\mathbf{p}}$. Since the number of terms in $\sigma_{\mathbf{p}}$ is, by Lemma 4.2, $O(\frac{1}{n}(P_{1}\cdots P_{h})^{\frac{1}{k}}+(P_{1}\cdots P_{h})^{\frac{1}{k}-\frac{1}{h}-\delta})$, we have:
$S_{1}$ $\leq\sum_{\mathbf{p}\in\mathcal{P}}\sigma_{\mathbf{p}}$
$\ll\sum_{\mathbf{p}\in\mathcal{P}}\bigg{(}\frac{1}{n}(P_{1}\cdots P_{h})^{\frac{1}{k}}+(P_{1}\cdots P_{h})^{\frac{1}{k}-\frac{1}{h}-\delta}\bigg{)}\,\frac{f(P_{1})}{P_{1}^{\frac{1}{k}}}\cdots\frac{f(P_{h})}{P_{h}^{\frac{1}{k}}}$
$=\frac{1}{n}\sum_{\mathbf{p}\in\mathcal{P}}f(P_{1})\cdots f(P_{h})+\sum_{\mathbf{p}\in\mathcal{P}}\frac{f(P_{1})}{P_{1}^{\frac{1}{h}+\delta}}\cdots\frac{f(P_{h})}{P_{h}^{\frac{1}{h}+\delta}}$
$\ll\frac{1}{n}\Bigg{(}\sum_{j=0}^{L}f(2^{j})\Bigg{)}\Bigg{(}\sum_{j=0}^{J}f(2^{j})\Bigg{)}^{h-1}+\Bigg{(}\sum_{j=0}^{L}\frac{f(2^{j})}{(2^{j})^{\frac{1}{h}+\delta}}\Bigg{)}\Bigg{(}\sum_{j=0}^{J}\frac{f(2^{j})}{(2^{j})^{\frac{1}{h}+\delta}}\Bigg{)}^{h-1}.$ (4.3)
Since $f(x)=x^{\omega}\phi(x)$, we have $\sum_{j=0}^{J}f(2^{j})\ll_{\varepsilon}f(n)\sum_{j=0}^{J}2^{-j(\omega-\varepsilon)}\ll f(n)$, we have
$\frac{1}{n}\Bigg{(}\sum_{j=0}^{L}f(2^{j})\Bigg{)}\Bigg{(}\sum_{j=0}^{J}f(2^{j})\Bigg{)}^{h-1}\ll\frac{f(n/\epsilon(n))f(n)^{h-1}}{n}=o\bigg{(}\frac{f(n)^{h}}{n}\bigg{)}.$
For the other term, if $\omega>1/h$ then
$\sum_{j=0}^{J}\frac{f(2^{j})}{(2^{j})^{\frac{1}{h}+\delta}}$ $\ll_{\varepsilon}\frac{f(n)}{n^{\frac{1}{h}+\delta}}\sum_{j=0}^{J}(2^{-j})^{\omega-(\frac{1}{h}+\delta)-\varepsilon}\ll\frac{f(n)}{n^{1/h}},$
and if $\omega=1/h$ then $\sum_{j=0}^{J}\frac{f(2^{j})}{(2^{j})^{\frac{1}{h}+\delta}}\ll_{\varepsilon}\sum_{j=0}^{J}2^{-j(\delta-\varepsilon)}\ll 1=o(f(n)^{h}/n)$. Therefore, it follows from (4.3) that $S_{1}=o(f(n)^{h}/n)$.
For $S_{2}$, since $f(x)=x^{\omega}\phi(x)$, by the definition of $\epsilon(x)$ we have
$S_{2}$ $\sim\phi(n)^{h}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{N}^{k}\\ x_{1}+\cdots+x_{h}=n\\ x_{j}\geq n/\epsilon(n),\,\forall j\end{subarray}}(x_{1}\cdots x_{h})^{\omega-\frac{1}{k}}\frac{\phi(x_{1})}{\phi(n)}\cdots\frac{\phi(x_{h})}{\phi(n)}$
$\sim\phi(n)^{h}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{N}^{k}\\ x_{1}+\cdots+x_{h}=n\\ x_{j}\geq n/\epsilon(n),\,\forall j\end{subarray}}(x_{1}\cdots x_{h})^{\omega-\frac{1}{k}}$ (4.4)
Using the same methods used to calculate $S_{1}$, one can show that
$S_{3}:=\phi(n)^{h}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{N}^{k}\\ x_{1}+\cdots+x_{h}=n\\ \exists j\,|\,x_{j}<n/\epsilon(n)\end{subarray}}(x_{1}\cdots x_{h})^{\omega-\frac{1}{k}}=o\bigg{(}\frac{f(n)^{h}}{n}\bigg{)}.$
Thus, from Lemma 4.1, (4.4) implies that
$S_{2}\sim\frac{1}{k^{h}}\frac{\Gamma(\omega)^{h}}{\Gamma(h\omega)}\mathfrak{S}_{k,h}(n)\frac{f(n)^{h}}{n},$
concluding the proof. ∎
Lemma 4.3 implies that $\mathbb{N}^{k}$ satisfies condition (iii) for every admissible regularly varying $F$ with $H_{\textnormal{(iii)}}\leq\max\{H_{V},H_{W}\}\ll 8^{k}k^{2}$. Therefore, since $\max\{2H_{\textnormal{(ii)}}+1,H_{\textnormal{(iii)}}\}\ll 8^{k}k^{2}$, Theorem 1.1 (ii) follows from Theorem 1.4.
5. Proof of Theorem 1.2
Fix $k\geq 1$ an integer, $h\geq 2^{k}+1$, let $\omega\geq 1/h$ be a real number, and $N$ be a large integer. For the next two sections, we will deal with $\mathbb{P}^{k}$. Define
$T(\alpha;x)=\sum_{n\leq x}n^{\omega}\frac{\mathbbm{1}_{\mathbb{P}^{k}}(n)}{n^{1/k}/\log n}\,e(\alpha n),\qquad\widetilde{T}(\alpha;x)=\sum_{x/h\leq n\leq x}n^{\omega}\frac{\mathbbm{1}_{\mathbb{P}^{k}}(n)}{n^{1/k}/\log n}\,e(\alpha n),$ (5.1)
so that, by the orthogonality of the $e(N\alpha)$,
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{P}^{k}\\ x_{1}+\cdots+x_{h}=N\end{subarray}}(x_{1}\cdots x_{h})^{\omega-\frac{1}{k}}(\log x_{1}\cdots\log x_{h})=\int_{0}^{1}T(\alpha;N)^{h}\,e(-N\alpha)\,\mathrm{d}\alpha.$
If $x_{1}+\cdots+x_{h}=N$, then $x_{i}\geq N/h$ for at least one $x_{i}$. Hence,
$\int_{0}^{1}(T(\alpha;N)-\widetilde{T}(\alpha;N))^{h}\,e(-N\alpha)\,\mathrm{d}\alpha=0.$
Therefore,
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{P}^{k}\\ x_{1}+\cdots+x_{h}=N\end{subarray}}(x_{1}\cdots x_{h}$ $)^{\omega-\frac{1}{k}}(\log x_{1}\cdots\log x_{h})$
$=\int_{0}^{1}\big{(}T(\alpha;N)^{h}-(T(\alpha;N)-\widetilde{T}(\alpha;N))^{h}\big{)}\,e(-N\alpha)\,\mathrm{d}\alpha$
$=\sum_{j=1}^{h}(-1)^{j+1}\binom{h}{j}\int_{0}^{1}\widetilde{T}(\alpha;N)^{j}\,T(\alpha;N)^{h-j}\,e(-N\alpha)\,\mathrm{d}\alpha.$ (5.2)
By (5.2), Theorem 1.2 will follow from a combination of Propositions 5.1 and 5.9 below.
5.1. Major arcs
Let $N$ be a large integer. As is standard in the circle method, we divide $[0,1]$ into major arcs $\mathfrak{M}$ and minor arcs $\mathfrak{m}$, where the major arcs $\mathfrak{M}=\mathfrak{M}_{N}$ are defined as
$\mathfrak{M}:=\bigsqcup_{q\leq Q}\bigsqcup_{\begin{subarray}{c}a=1\\ (a,q)=1\end{subarray}}^{q}\mathfrak{M}(q,a),\quad\mathfrak{M}(q,a):=\bigg{\{}\alpha\in[0,1]\leavevmode\nobreak\ \bigg{|}\leavevmode\nobreak\ \bigg{|}\alpha-\frac{a}{q}\bigg{|}\leq\frac{Q}{N}\bigg{\}}$
for a parameter $Q=Q(N):=(\log N)^{C}$, where $C>1$ is a large fixed constant. The minor arcs are defined as the complement, $\mathfrak{m}=\mathfrak{m}_{N}:=[0,1]\setminus\mathfrak{M}$. Splitting the integral in (5.2) as $\int_{0}^{1}=\int_{\mathfrak{M}}+\int_{\mathfrak{m}}$, we anticipate that the contribution from the minor arcs is negligible (see subsection 5.2). The goal of this subsection is to prove the following:
Proposition 5.1 (Major arcs).
We have
$\sum_{j=1}^{h}(-1)^{j+1}\binom{h}{j}\int_{\mathfrak{M}}\widetilde{T}(\alpha;N)^{j}\,T(\alpha;\,N)^{h-j}\,$ $e(-N\alpha)\,\mathrm{d}\alpha$
$=\mathfrak{S}^{*}_{k,h}(N)\,\frac{\Gamma(\omega)^{h}}{\Gamma(h\omega)}\,N^{h\omega-1}+O_{R}\bigg{(}\frac{N^{h\omega-1}}{(\log N)^{R}}\bigg{)},$
where $\mathfrak{S}^{*}_{k,h}(N)$ is as in (1.3), and $R(C)\to\infty$ as $C\to\infty$.
5.1.1. Lemmas
Given $q\geq 1$, $1\leq a\leq q$, define
$\pi(x;q,a):=\sum_{\begin{subarray}{c}n\leq x\\ n\equiv a(q)\end{subarray}}\mathbbm{1}_{\mathbb{P}}(n).$
We start with the classical estimate for primes in arithmetic progressions due to Siegel and Walfisz (cf. Montgomery–Vaughan [11, Corollary 11.21]).
Lemma 5.2 (Siegel–Walfisz).
Uniformly for $q\leq Q$, $1\leq a\leq q$ with $(a,q)=1$, we have
$\pi(x;q,a)=\bigg{(}1+O_{C,R}\bigg{(}\frac{1}{(\log x)^{R}}\bigg{)}\bigg{)}\frac{\mathrm{li}(x)}{\varphi(q)}$
for every $R>1$, where $\mathrm{li}(x)=\int_{2^{-}}^{x}(\log t)^{-1}\,\mathrm{d}t$.
Corollary 5.3.
Uniformly for $q\leq Q$, $1\leq a\leq q$ with $(a,q)=1$, we have
$\sum_{\begin{subarray}{c}n\leq x\\ n\equiv a(q)\end{subarray}}n^{\omega}\frac{\mathbbm{1}_{\mathbb{P}^{k}}(n)}{n^{1/k}/\log n}=\bigg{(}1+O_{C,R}\bigg{(}\frac{1}{(\log x)^{R}}\bigg{)}\bigg{)}\frac{P_{k}(q,a)}{\varphi(q)}\frac{x^{\omega}}{\omega},$
for every $R>1$, where $P_{k}(q,a):=|\{1\leq r\leq q\leavevmode\nobreak\ |\leavevmode\nobreak\ r^{k}\equiv a\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ q)\}|$.
Proof.
Write $P=P_{k}(q,a)$, and let $1\leq r_{1},\ldots,r_{P}\leq q$ be such that $r_{i}^{k}\equiv a\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ q)$. Since $\{n\leq x\leavevmode\nobreak\ |\leavevmode\nobreak\ n\in\mathbb{P}^{k},n\equiv a\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ q)\}=\sum_{i=1}^{P}\pi(x^{1/k};q,r_{i})$, we have
$\sum_{\begin{subarray}{c}n\leq x\\ n\equiv a(q)\end{subarray}}n^{\omega}\frac{\mathbbm{1}_{\mathbb{P}^{k}}(n)}{n^{1/k}/\log n}$ $=\sum_{i=1}^{P}\int_{2^{-}}^{x}\frac{t^{\omega}}{t^{1/k}/\log t}\,\mathrm{d}\pi(t^{1/k};q,r_{i})$
$=\frac{P}{\varphi(q)}\frac{x^{\omega}}{\omega}+\int_{2^{-}}^{x}\frac{t^{\omega}}{t^{1/k}/\log t}\,\mathrm{d}\bigg{(}\pi(t^{1/k};q,r_{i})-\frac{\mathrm{li}(t^{1/k})}{\varphi(q)}\bigg{)}.$
By partial summation, Lemma 5.2 implies that
$\int_{2^{-}}^{x}\frac{t^{\omega}}{t^{1/k}/\log t}\,$ $\mathrm{d}\bigg{(}\pi(t^{1/k};q,r_{i})-\frac{\mathrm{li}(t^{1/k})}{\varphi(q)}\bigg{)}$
$\hskip 30.00005pt=\bigg{(}\pi(x^{1/k};q,r_{i})-\frac{\mathrm{li}(x^{1/k})}{\varphi(q)}\bigg{)}\frac{x^{\omega}}{x^{1/k}/\log x}$
$\hskip 30.00005pt\phantom{=}\ -\int_{2^{-}}^{x}\bigg{(}\pi(t^{1/k};q,r_{i})-\frac{\mathrm{li}(t^{1/k})}{\varphi(q)}\bigg{)}\,t^{\omega-\frac{1}{k}-1}(1+(\omega-\tfrac{1}{k})\log t)\,\mathrm{d}t$
$\hskip 30.00005pt\ll_{C,R}\frac{x^{\omega}}{(\log x)^{R}},$
which, since $P\leq k$, concludes the proof. ∎
Next, we prove a lemma ensures that the singular series $\mathfrak{S}_{k,h}^{*}(N)$ converges absolutely and is bounded away from 00 and $\infty$ for $N\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$, where $K(k)$ is as in (1.2).
Lemma 5.4 (Singular series).
For $N\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$, we have $\mathfrak{S}_{k,h}^{*}(N)\asymp 1$. Moreover, let
$\mathfrak{S}_{k,h}^{*}(N,Q):=\sum_{q\leq Q}\sum_{\begin{subarray}{c}a=1\\ (a,q)=1\end{subarray}}^{q}\frac{S(q,a)^{h}}{\varphi(q)^{h}}e\bigg{(}-\frac{Na}{q}\bigg{)},$
where $S(q,a)$ be as in (1.1). Then, for every $\varepsilon>0$,
$\mathfrak{S}_{k,h}^{*}(N,Q)=\bigg{(}1+O_{\varepsilon}\bigg{(}\frac{1}{Q^{1/2^{k-1}-\varepsilon}}\bigg{)}\bigg{)}\mathfrak{S}_{k,h}^{*}(N)\quad\text{as}\quad Q\to\infty.$
Proof.
By Lemmas 8.10 and 8.12 in Hua [8], we have $\mathfrak{S}_{k,\ell}^{*}(N)\gg 1$ for $N\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$ whenever $\ell\geq 3$ for $k=1$, $\ell\geq 5$ for $k=2$, and $\ell\geq 3k$ for $k\geq 3$. In all cases, this holds true for $h\geq 2^{k}+1$.
For the second part, let $q\geq 1$, $1\leq a\leq q$ with $(a,q)=1$. By Weyl’s inequality we have $S(q,a)\ll_{\varepsilon}q^{1-1/2^{k-1}+\varepsilon}$ uniformly in $q$ (cf. Nathanson [12, Theorem 4.4]). Thus, since $\varphi(q)\gg_{\varepsilon}q^{1-\varepsilon}$ and $h\geq 2^{k}+1$, we have
$|\mathfrak{S}_{k,h}^{*}(N)-\mathfrak{S}_{k,h}^{*}(N,Q)|$ $\leq\sum_{q>Q}\sum_{\begin{subarray}{c}a=1\\ (a,q)=1\end{subarray}}^{q}\bigg{|}\frac{S(q,a)^{h}}{\varphi(q)^{h}}\bigg{|}$
$\ll_{\varepsilon}\sum_{q>Q}(q^{-1/2^{k-1}+\varepsilon})^{h}\,q$
$\ll_{\varepsilon}\sum_{q>Q}q^{-1-1/2^{k-1}+\varepsilon}\asymp Q^{-1/2^{k-1}+\varepsilon}.\qed$
5.1.2. Integral over major arcs
In order to calculate the integral in Proposition 5.1, we will approximate $T$ and $\widetilde{T}$ in terms of, respectively,
$u(\vartheta;x):=\sum_{n\leq x}n^{\omega-1}e(n\vartheta),\qquad\widetilde{u}(\vartheta;x):=\sum_{n\leq x}n^{\omega-1}e(n\vartheta).$ (5.3)
Lemma 5.5.
Let $q\leq Q$, and $1\leq a\leq q$ with $(a,q)=1$. For $\alpha\in\mathfrak{M}(q,a)$, write $\vartheta:=\alpha-\frac{a}{q}$. Then, for every $R>1$, we have
$T(\alpha;N)=\frac{S(q,a)}{\varphi(q)}u(\vartheta;N)+O_{R}\bigg{(}\frac{N^{\omega}}{(\log N)^{R}}\bigg{)},$
$\widetilde{T}(\alpha;N)=\frac{S(q,a)}{\varphi(q)}\widetilde{u}(\vartheta;N)+O_{R}\bigg{(}\frac{N^{\omega}}{(\log N)^{R}}\bigg{)}$
uniformly in $q\leq Q$.
Proof.
We prove the lemma for $T$; the argument for $\widetilde{T}$ is analogous. By Lemma 5.3, we have
$T\bigg{(}\frac{a}{q};x\bigg{)}$ $=\sum_{n\leq x}n^{\omega}\frac{\mathbbm{1}_{\mathbb{P}^{k}}(n)}{n^{1/k}/\log n}e\bigg{(}\frac{na}{q}\bigg{)}$
$=\sum_{\begin{subarray}{c}r=1\\ (r,q)=1\end{subarray}}^{q}\Bigg{(}\sum_{\begin{subarray}{c}n\leq x\\ n\equiv r(q)\end{subarray}}n^{\omega}\frac{\mathbbm{1}_{\mathbb{P}^{k}}(n)}{n^{1/k}/\log n}\Bigg{)}e\bigg{(}\frac{ra}{q}\bigg{)}+O(Q)$
$=\Bigg{(}\sum_{\begin{subarray}{c}r=1\\ (r.q)=1\end{subarray}}^{q}P_{k}(q,r)\,e\bigg{(}\frac{ra}{q}\bigg{)}\Bigg{)}\frac{1}{\varphi(q)}\frac{x^{\omega}}{\omega}+O_{R}\bigg{(}\frac{x^{\omega}}{(\log x)^{R}}\bigg{)}$
$=\frac{S(q,a)}{\varphi(q)}\frac{x^{\omega}}{\omega}+O_{R}\bigg{(}\frac{x^{\omega}}{(\log x)^{R}}\bigg{)}.$ (5.4)
(Note that $\widetilde{T}(\frac{a}{q};x)=T(\frac{a}{q};x)-T(\frac{a}{q};\frac{x}{h}-1)$.) We have
$T(\alpha;N)-\frac{S(q,a)}{\varphi(q)}u(\vartheta;N)=\sum_{n\leq N}\underbrace{\Bigg{(}n^{\omega}\frac{\mathbbm{1}_{\mathbb{P}}(n)}{n^{1/k}/\log n}e\bigg{(}\frac{na}{q}\bigg{)}-\frac{S(q,a)}{\varphi(q)}n^{\omega-1}\Bigg{)}}_{=:\,D(n)}e(n\vartheta)$
By (5.4), and the fact that $\sum_{n\leq t}n^{\omega-1}=\frac{t^{\omega}}{\omega}+O(1)$, we have $\sum_{n\leq t}D(n)=O_{A}(t/(\log t)^{A})$. By using partial summation we conclude, since $|\vartheta|\leq Q/N=(\log N)^{C}/N$, that
$T(\alpha;N)-\frac{S(q,a)}{\varphi(q)}u(\vartheta;N)$ $=\bigg{(}\sum_{n\leq N}D(n)\bigg{)}e(N\vartheta)-2\pi i\vartheta\int_{1}^{N}\bigg{(}\sum_{n\leq t}D(n)\bigg{)}e(t\vartheta)\,\mathrm{d}t$
$\ll\bigg{|}\sum_{n\leq N}D(n)\bigg{|}+N|\vartheta|\max_{t\leq N}\bigg{(}\sum_{n\leq t}D(n)\bigg{)}$
$=O_{R}\bigg{(}\frac{N^{\omega}}{(\log N)^{R}}\bigg{)}.\qed$
Lemma 5.6.
For $-\frac{1}{2}\leq\vartheta\leq\frac{1}{2}$, $\vartheta\neq 0$, we have
$|u(\vartheta;N)|\ll\min\{N^{\omega},|\vartheta|^{-\omega}\},\qquad|\widetilde{u}(\vartheta;N)|\ll N^{\omega-1}|\vartheta|^{-1},$
Proof.
We have $|u(\vartheta;N)|\leq u(0;N)=\frac{1}{\omega}N^{\omega}+O(1)$. On the other hand, let $M:=\lfloor|\vartheta|^{-1}\rfloor$, so that
$|u(\vartheta;N)|\leq\sum_{n\leq M}n^{\omega-1}+\bigg{|}\sum_{n=M+1}^{N}n^{\omega-1}e(n\vartheta)\bigg{|}\ll M^{\omega}+\bigg{|}\sum_{n=M+1}^{N}n^{\omega-1}e(n\vartheta)\bigg{|}.$ (5.5)
Let $E(x):=\sum_{n\leq x}e(n\vartheta)$. Since $|1-e(\vartheta)|=|e(-\vartheta/2)-e(\vartheta/2)|=2|\sin\pi\vartheta|\geq 2|\vartheta|$ for $\vartheta$ in the range considered, we have
$|E(x)|=\bigg{|}\sum_{n\leq x}e(n\vartheta)\bigg{|}=\bigg{|}\frac{1-e((\lfloor x\rfloor+1)\vartheta)}{1-e(\vartheta)}\bigg{|}\leq\frac{1}{|\vartheta|^{-1}}.$
By using that $E(x)\ll|\vartheta|^{-1}$, we obtain from partial summation that
$\sum_{n=M+1}^{N}n^{\omega-1}e(n\vartheta)$ $=N^{\omega-1}E(N)-M^{\omega-1}E(M)-(\omega-1)\int_{M}^{N}E(t)t^{\omega-2}\,\mathrm{d}t$
$\ll(N^{\omega-1}+M^{\omega-1})\,|\vartheta|^{-1}\ll M^{\omega}\asymp|\vartheta|^{-\omega},$
so the result follows from (5.5).
Similarly, for $\widetilde{u}$, we have
$\widetilde{u}(\vartheta;N)$ $=\sum_{N/h\leq n\leq N}n^{\omega-1}e(n\vartheta)$
$=N^{\omega-1}E(N)-\bigg{(}\frac{N}{h}\bigg{)}^{\omega-1}E(N/h)-(\omega-1)\int_{N/h}^{N}E(t)t^{\omega-2}\,\mathrm{d}t$
$\ll N^{\omega-1}|\vartheta|^{-1}.\qed$
Lemma 5.7.
For $1\leq j\leq h$, we have, for every $\varepsilon>0$,
$\int_{\mathfrak{M}}\widetilde{T}(\alpha;N)^{j}\,$ $T(\alpha;N)^{h-j}\,e(-N\alpha)\,\mathrm{d}\alpha$
$=\bigg{(}1+O_{\varepsilon}\bigg{(}\frac{1}{(\log N)^{C/2^{k-1}-\varepsilon}}\bigg{)}\bigg{)}\mathfrak{S}^{*}_{k,h}(N)\int_{0}^{1}\widetilde{u}(\vartheta;N)^{j}\,u(\vartheta;N)^{h-j}\,e(-N\vartheta)\,\mathrm{d}\vartheta$
$\hskip 225.00034pt+O\bigg{(}\frac{N^{h\omega-1}}{(\log N)^{C(h-1)\omega}}\bigg{)}.$
Proof.
Since $|u(\vartheta;N)|\leq u(0;N)=\frac{1}{\omega}N^{\omega}+O(1)$, Lemma 5.5 implies that
$T(\alpha;N)^{h}=\frac{S(q,a)^{h}}{\varphi(q)^{h}}u(\vartheta;N)^{h}+O_{R}\bigg{(}\frac{N^{h\omega}}{(\log N)^{R}}\bigg{)},$
and similarly for $\widetilde{T}$, $\widetilde{u}$. Using that $\int_{\mathfrak{M}(q,a)}\mathrm{d}\alpha=2Q/N$, this leads to the estimate
$\int_{\mathfrak{M}}$ $\widetilde{T}(\alpha;N)^{j}\,T(\alpha;N)^{h-j}\,e(-N\alpha)\,\mathrm{d}\alpha$
$=\sum_{q\leq Q}\sum_{\begin{subarray}{c}a=1\\ (a,q)=1\end{subarray}}^{q}\int_{\mathfrak{M}(q,a)}\widetilde{T}(\alpha;N)^{j}\,T(\alpha;N)^{h-j}\,e(-N\alpha)\,\mathrm{d}\alpha$
$=\sum_{q\leq Q}\sum_{\begin{subarray}{c}a=1\\ (a,q)=1\end{subarray}}^{q}\frac{S(q,a)^{h}}{\varphi(q)^{h}}\int_{\mathfrak{M}(q,a)}\widetilde{u}\bigg{(}\alpha-\frac{a}{q};N\bigg{)}^{j}\,u\bigg{(}\alpha-\frac{a}{q};N\bigg{)}^{h-j}e(-N\alpha)\,\mathrm{d}\alpha$
$\hskip 270.00041pt+O_{R}\bigg{(}\frac{QN^{\omega h-1}}{(\log N)^{R}}\bigg{)}$
$=\underbrace{\sum_{q\leq Q}\sum_{\begin{subarray}{c}a=1\\ (a,q)=1\end{subarray}}^{q}\frac{S(q,a)^{h}}{\varphi(q)^{h}}e\bigg{(}-\frac{Na}{q}\bigg{)}}_{=\,\mathfrak{S}_{h}(N,Q)}\int_{-Q/N}^{Q/N}\widetilde{u}(\vartheta;N)^{j}u(\vartheta;N)^{h-j}\,e(-N\vartheta)\,\mathrm{d}\vartheta$
$\hskip 270.00041pt+O_{R}\bigg{(}\frac{N^{\omega h-1}}{(\log N)^{R}}\bigg{)}$
$=\bigg{(}1+O_{\varepsilon}\bigg{(}\frac{1}{(\log N)^{C/2^{k-1}-\varepsilon}}\bigg{)}\bigg{)}\mathfrak{S}^{*}_{k,h}(N)\int_{-Q/N}^{Q/N}\widetilde{u}(\vartheta;N)^{j}u(\vartheta;N)^{h-j}\,e(-N\vartheta)\,\mathrm{d}\vartheta$
$\hskip 270.00041pt+O_{R}\bigg{(}\frac{N^{\omega h-1}}{(\log N)^{R}}\bigg{)},$
where the last line follows from Lemma 5.4. By Lemma 5.6, we have
$\bigg{|}\int_{[-\frac{1}{2},-\frac{Q}{N}]\cup[\frac{Q}{N},\frac{1}{2}]}\widetilde{u}(\vartheta;N)^{j}u(\vartheta;N)^{h-j}e(-N\vartheta)\,\mathrm{d}\vartheta\bigg{|}$ $\leq\int_{[-\frac{1}{2},-\frac{Q}{N}]\cup[\frac{Q}{N},\frac{1}{2}]}|\widetilde{u}(\vartheta;N)|^{j}\,|u(\vartheta;N)|^{h-j}\,\mathrm{d}\vartheta$
$\ll\frac{N^{j\omega}}{Q^{j}}\int_{Q/N}^{1/2}\vartheta^{-(h-j)\omega}\,\mathrm{d}\vartheta$
$\ll\frac{N^{h\omega-1}}{Q^{h\omega-1+j(1-\omega)}}\ll\frac{N^{h\omega-1}}{Q^{(h-1)\omega}}.$
Since $\mathfrak{S}_{k,h}^{*}(N)\asymp 1$, this finishes the proof. ∎
5.1.3. The singular integral
Lemma 5.7 reduces Proposition 5.1 to the calculation of the integral in $u$, $\widetilde{u}$ that appears in its statement, which is done in the following lemma:
Lemma 5.8.
We have
$\sum_{j=1}^{h}(-1)^{j+1}\binom{h}{j}\int_{0}^{1}\widetilde{u}(\vartheta;N)^{j}\,u(\vartheta;N)^{h-j}\,e(-N\vartheta)\,\mathrm{d}\vartheta=(1+O(N^{-\omega}))\frac{\Gamma(\omega)^{h}}{\Gamma(h\omega)}N^{h\omega-1}$
Proof.
By the same argument used to obtain (5.2), we deduce from (5.3) that
$\sum_{j=1}^{h}(-1)^{j+1}\binom{h}{j}\int_{0}^{1}\widetilde{u}(\vartheta;N)^{j}\,u(\vartheta;N)^{h-j}\,e(-N\vartheta)\,\mathrm{d}\vartheta$ $=\int_{0}^{1}u(\vartheta;N)^{h}\,e(-N\vartheta)\,\mathrm{d}\vartheta$
$=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{Z}_{\geq 1}\\ x_{1}+\cdots+x_{h}=N\end{subarray}}(x_{1}\cdots x_{h})^{\omega-1},$
For $\ell\geq 2$ and $n\in\mathbb{Z}_{\geq 1}$, we claim that
$S_{\ell}(n):=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in\mathbb{Z}_{\geq 1}\\ x_{1}+\cdots+x_{\ell}=n\end{subarray}}(x_{1}\cdots x_{\ell})^{\omega-1}=\frac{\Gamma(\omega)^{\ell}}{\Gamma(\ell\omega)}n^{\ell\omega-1}+O(n^{(\ell-1)\omega-1}),$ (5.6)
which will finish the proof. The case $\ell=2$ is a consequence of Lemma 2.9 of Vaughan [15], which says that
$\sum_{m=1}^{n-1}m^{a-1}(n-m)^{b-1}=\frac{\Gamma(a)\Gamma(b)}{\Gamma(a+b)}n^{a+b-1}+O_{a,b}(n^{a-1})$ (5.7)
for any real numbers $a\geq b>0$. Assuming by induction that (5.6) holds for some $\ell\geq 2$, we obtain from (5.7) that
$S_{\ell+1}(n)$ $=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{\ell}\in\mathbb{Z}_{\geq 1}\\ x_{1}+\cdots+x_{\ell}\leq n\end{subarray}}(x_{1}\cdots x_{\ell})^{\omega-1}\big{(}n-(x_{1}+\cdots+x_{\ell})\big{)}^{\omega-1}$
$=\sum_{m=1}^{n-1}S_{\ell}(m)\,(n-m)^{\omega-1}$
$=\frac{\Gamma(\omega)^{\ell}}{\Gamma(\ell\omega)}\sum_{m=1}^{n-1}m^{\ell\omega-1}(n-m)^{\omega-1}+O\bigg{(}\sum_{m=1}^{n-1}m^{(\ell-1)\omega-1}(n-m)^{\omega-1}\bigg{)}$
$=\frac{\Gamma(\omega)^{\ell+1}}{\Gamma((\ell+1)\omega)}n^{(\ell+1)\omega-1}+O(n^{\ell\omega-1}),$
concluding the proof. ∎
Proof of Proposition 5.1.
Since both $C/2^{k-1}-\varepsilon\to\infty$ and $C(h-1)\omega\to\infty$ as $C\to\infty$, by Lemmas 5.7 and 5.8, we have
$\sum_{j=1}^{h}(-1)^{j+1}\binom{h}{j}\int_{\mathfrak{M}}\widetilde{T}(\alpha;N)^{j}\,T(\alpha;N)^{h-j}\,e(-N\alpha)\,\mathrm{d}\alpha$
$=\bigg{(}1+O_{R}\bigg{(}\frac{1}{(\log N)^{R}}\bigg{)}\bigg{)}\mathfrak{S}^{*}_{k,h}(N)\sum_{j=1}^{h}(-1)^{j+1}\binom{h}{j}\int_{0}^{1}\widetilde{u}(\vartheta;N)^{j}\,u(\vartheta;N)^{h-j}\,e(-N\vartheta)\,\mathrm{d}\vartheta$
$\hskip 295.00044pt+O\bigg{(}\frac{N^{h\omega-1}}{(\log N)^{R}}\bigg{)}$
$=\bigg{(}1+O\bigg{(}\frac{1}{(\log N)^{R}}\bigg{)}\bigg{)}(1+O(N^{-\omega}))\mathfrak{S}^{*}_{k,h}(N)\frac{\Gamma(\omega)^{h}}{\Gamma(h\omega)}N^{h\omega-1}+O\bigg{(}\frac{N^{h\omega-1}}{(\log N)^{R}}\bigg{)},$
where $R=R(C)\to\infty$ as $C\to\infty$. Since $\mathfrak{S}^{*}_{k,h}(N)\ll 1$ by Lemma 5.4, rearranging the error terms concludes the proof. ∎
5.2. Minor arcs
We now turn our attention to the minor arcs $\mathfrak{m}$, as defined in the beginning of the last subsection. Together with (5.2) and Proposition 5.1, the next result directly implies Theorem 1.2.
Proposition 5.9 (Minor arcs).
We have
$\bigg{|}\sum_{j=1}^{h}\int_{\mathfrak{m}}\widetilde{T}(\alpha;N)^{j}\,T(\alpha;N)^{h-j}\,e(-N\alpha)\,\mathrm{d}\alpha\bigg{|}\ll_{R}\frac{N^{h\omega-1}}{(\log N)^{R}},$
where $R=R(C)\to\infty$ as $C\to\infty$.
5.2.1. Lemmas
Define the function
$g(\alpha;x):=\sum_{n\leq x}\mathbbm{1}_{\mathbb{P}^{k}}(n)e(n\alpha).$
The first lemma we need is equivalent to showing that $\mathbb{P}^{k}$ satisfies condition (ii) of Theorem 1.4 with $H_{\textnormal{(ii)}}\leq 2^{k-1}$.
Lemma 5.10 (Mean-value estimate).
There exists $c(k)\in\mathbb{R}$ such that, for every $\ell\geq 2^{k-1}$,
$\sum_{n\leq x}r_{\mathbb{P}^{k},\ell}(n)^{2}\leq x^{2\ell/k-1}(\log x)^{c(k)}.$
Proof.
By Parseval’s identity, we have
$\sum_{n\leq x}r_{\mathbb{P}^{k},2^{k-1}}(n)^{2}\leq\int_{0}^{1}|g(\alpha;x)|^{2^{k}}\,\mathrm{d}\alpha\leq\sum_{n\leq 2^{k-1}x}r_{\mathbb{P}^{k},2^{k-1}}(n)^{2},$
where the middle term, by the orthogonality of the $e(\alpha n)$, counts the number of solutions to the equation
$n_{1}+\cdots+n_{2^{k-1}}=m_{1}+\cdots+m_{2^{k-1}}$ (5.8)
with $n_{i},m_{j}\leq x$ and $n_{i},m_{j}\in\mathbb{P}^{k}$. By Hua’s inequality (cf. Hua [8, Theorem 4]), there exists $c(k)\in\mathbb{R}$ such that
$\int_{0}^{1}\bigg{|}\sum_{1\leq n\leq x}e(\alpha n^{k})\bigg{|}^{2\ell}\,\mathrm{d}\alpha\ll x^{2\ell-k}(\log x)^{c(k)}.$
for every $\ell\geq 2^{k-1}$. In other words, the number of solutions to (5.8) with $n_{i},m_{j}\leq x$, $n_{i},m_{j}\in\mathbb{N}^{k}$ is $O(x^{2^{k}/k-1}(\log x)^{c(k)})$, which is an upper bound to the number of solutions with $n_{i},m_{j}\in\mathbb{P}^{k}$. Therefore, since $|g(\alpha;x)|\leq g(0;x)\asymp x^{1/k}/\log x$, for $\ell\geq 2^{k-1}$ we have
$\int_{0}^{1}|g(\alpha;x)|^{2\ell}\,\mathrm{d}\alpha$ $\leq|g(0;x)|^{2\ell-2^{k}}\,\int_{0}^{1}|g(\alpha;x)|^{2^{k}}\,\mathrm{d}\alpha$
$\ll\bigg{(}\frac{x^{1/k}}{\log x}\bigg{)}^{2\ell-2^{k}}\,x^{2^{k}/k-1}(\log x)^{c(k)}$
$\ll x^{2\ell/k-1}(\log x)^{c(k)}.\qed$
Next, we bound the value of $g(\alpha;x)$ for $\alpha\in\mathfrak{m}$. To this end, we employ a lemma due to Harman, in the form stated by Kumchev–Tolev [10, Lemma 3.3].
Lemma 5.11 (Harman [7]).
Let $\alpha\in[0,1]$ be such that $|\alpha-\frac{a}{q}|\leq q^{-2}$, where $a$, $q$ are integers with $1\leq q\leq x$ and $(a,q)=1$. Then, there exists a constant $G=G(k)>0$ such that
$g(\alpha;x)\ll\big{(}q^{-1}+x^{-1/2k}+q^{1/2}x^{-1}\big{)}^{4^{1-k}}x^{1/k}(\log x)^{G}$
Corollary 5.12.
$\sup_{\alpha\in\mathfrak{m}}|g(\alpha;N)|\ll\frac{N^{1/k}}{(\log N)^{4^{1-k}C-G}}$.
Proof.
By Dirichlet’s theorem (cf. Nathanson [12, Theorem 4.1]), for any $\alpha\in[0,1]$ there exists $1\leq q\leq N/Q$ and $0\leq a\leq q$ such that
$\bigg{|}\alpha-\frac{a}{q}\bigg{|}\leq\frac{Q}{qN}\leq\min\bigg{(}\frac{Q}{N},\frac{1}{q^{2}}\bigg{)}.$
Since $\alpha\in\mathfrak{m}$, we must have $q>Q$ (otherwise, $\alpha\in\mathfrak{M}_{N}(q,a)$ by definition), so $Q<q\leq N/Q$. Hence, by Lemma 5.11,
$g(\alpha;N)$ $\ll\bigg{(}Q^{-1}+N^{-1/2k}+\bigg{(}\frac{N}{Q}\bigg{)}^{1/2}N^{-1}\bigg{)}^{4^{1-k}}N^{1/k}(\log N)^{G}$
$\ll\frac{N^{1/k}}{Q(\log N)^{-G}}=\frac{N^{1/k}}{(\log N)^{4^{1-k}C-G}}.\qed$
5.2.2. Integral over minor arcs
In order to prove Proposition 5.9, we first need two auxiliary lemmas concerning the integral of $\widetilde{T}$.
Lemma 5.13.
For $\ell\geq 2^{k}$, we have
$\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{\ell}\,\mathrm{d}\alpha\ll x^{\ell\omega-1}(\log x)^{c(k)+2^{k}}.$
Proof.
The integral $\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{2^{k}}\,\mathrm{d}\alpha$ counts the number of solutions to the Diophantine equation (5.8) with weights
$(n_{1}\ldots n_{2^{k-1}}m_{1}\cdots m_{2^{k-1}})^{\omega-\frac{1}{k}}(\log n_{1})\cdots(\log n_{2^{k-1}})(\log m_{1})\cdots$ $(\log m_{2^{k-1}})$
$\hskip 30.00005pt\ll\bigg{(}\frac{x^{\omega}}{x^{1/k}/\log x}\bigg{)}^{2^{k}},$
for $x/h\leq n_{i},m_{j}\leq x$. Thus, by considering this underlying equation, Lemma 5.10 yields the upper bound
$\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{2^{k}}\,\mathrm{d\alpha}\ll\frac{x^{2^{k}\omega}}{(x^{1/k}/\log x)^{2^{k}}}\int_{0}^{1}|g(\alpha;x)|^{2^{k}}\,\mathrm{d}\alpha\ll x^{2^{k}\omega-1}(\log x)^{c(k)+2^{k}}.$
Using the trivial estimate
$|\widetilde{T}(\alpha;x)|\leq\widetilde{T}(0;x)\ll x^{\omega-\frac{1}{k}}(\log x)\sum_{x/2\leq n\leq x}\mathbbm{1}_{\mathbb{P}^{k}}(n)\ll x^{\omega},$
for $\ell\geq 2^{k}$ we get
$\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{\ell}\,\mathrm{d}\alpha\ll x^{(\ell-2^{k})\omega}\int_{0}^{1}|\widetilde{T}(\alpha;x)|^{2^{k}}\,\mathrm{d}\alpha\ll x^{\ell\omega-1}(\log x)^{c(k)+2^{k}}.\qed$
Lemma 5.14.
We have
$\int_{\mathfrak{m}}|\widetilde{T}(\alpha;N)|^{h}\,\mathrm{d}\alpha\ll\frac{N^{\omega}}{(\log N)^{r(C)}},$
where $r(C)\to\infty$ as $C\to\infty$.
Proof.
By Corollary 5.12, for $\alpha\in\mathfrak{m}$, partial summation yields
$\widetilde{T}$ $(\alpha;N)$
$=\frac{N^{\omega}}{N^{1/k}/\log N}g(\alpha;N)-\frac{(N/h)^{\omega}}{(N/h)^{1/k}/(\log N/h)}g(\alpha;N/h)-\int_{N/h}^{N}g(\alpha;t)\,\mathrm{d}\bigg{(}\frac{t^{\omega}}{t^{1/k}/\log t}\bigg{)}$
$\ll N^{\omega-\frac{1}{k}}(\log N)\,\frac{N^{1/k}}{(\log N)^{4^{1-k}C-G}}=\frac{N^{\omega}}{(\log N)^{4^{1-k}C-G-1}}.$
Since $h\geq 2^{k}+1$, it follows from Lemma 5.13 that
$\int_{\mathfrak{m}}|\widetilde{T}(\alpha;N)|^{h}\,\mathrm{d}\alpha$ $\leq\bigg{(}\sup_{\alpha\in\mathfrak{m}}|\widetilde{T}(\alpha;N)|^{h-2^{k}}\bigg{)}\,\int_{0}^{1}|\widetilde{T}(\alpha;N)|^{2^{k}}\,\mathrm{d}\alpha$
$\ll\bigg{(}\frac{N^{\omega}}{(\log N)^{4^{1-k}C-G-1}}\bigg{)}^{h-2^{k}}N^{2^{k}\omega-1}(\log N)^{c(k)+2^{k}}$
$=\frac{N^{h\omega-1}}{(\log N)^{r(C)}},$
where $r(C)=(h-2^{k})(4^{1-k}C-G-1)-c(k)-2^{k}$. ∎
Proof of Proposition 5.9.
The series $T(\alpha;N)$ can be written as
$T(\alpha;N)=\sum_{j\leq\frac{\log N}{\log h}}\widetilde{T}^{\prime}(\alpha;h^{-j}N).$
where $\widetilde{T}^{\prime}(\alpha;x):=\sum_{x/h<n\leq x}n^{\omega}\frac{\mathbbm{1}_{\mathbb{P}^{k}}(n)}{n^{1/k}/\log n}\,e(\alpha n)$. Clearly, both Lemmas 5.13 and 5.14 apply to $\widetilde{T}^{\prime}$ as well. Therefore, by Hölder’s inequality [^4] With “$p=h$”, “$q=h/(h-1)$”. and Lemma 5.13, it follows that
$\int_{0}^{1}|T(\alpha;N)|^{h}\,\mathrm{d}\alpha$ $\leq(\log N)^{h-1}\int_{0}^{1}\bigg{(}\sum_{j\leq\frac{\log N}{\log h}}|\widetilde{T}^{\prime}(\alpha;h^{-j}N)|^{h}\bigg{)}\,\mathrm{d}\alpha$
$\ll(\log N)^{h}\max_{j\leq\frac{\log N}{\log h}}\int_{\mathfrak{m}}|\widetilde{T}^{\prime}(\alpha;h^{-j}N)|^{h}\,\mathrm{d}\alpha$
$\ll(\log N)^{h+c(k)+2^{k}}\,N^{h\omega-1}.$
By Lemma 5.14, another application of Hölder’s inequality yields
$\bigg{|}\sum_{j=1}^{h}\int_{\mathfrak{m}}\widetilde{T}(\alpha;N)^{j}\,T(\alpha;N)^{h-j}\,$ $e(-N\alpha)\,\mathrm{d}\alpha\bigg{|}$
$\leq\sum_{j=1}^{h}\bigg{(}\int_{\mathfrak{m}}|\widetilde{T}(\alpha;N)|^{h}\,\mathrm{d}\alpha\bigg{)}^{j/h}\,\bigg{(}\int_{0}^{1}|T(\alpha;N)|^{h}\,\mathrm{d}\alpha\bigg{)}^{1-j/h}$
$\leq\sum_{j=1}^{h}\bigg{(}\frac{N^{h\omega-1}}{(\log N)^{r(C)}}\bigg{)}^{j/h}\,\big{(}(\log N)^{h+c(k)+2^{k}}N^{h\omega-1}\big{)}^{1-j/h}$
$\ll\frac{N^{h\omega-1}}{(\log N)^{r(C)/h-(h+c(k)+2^{k})(h-1)/h}}.$
Writing $R=R(C):=r(C)/h-(h+c(k)+2^{k})(h-1)/h$, we see that choosing $C$ large, $R$ can be taken to be as large as desired, concluding the proof. ∎
6. Waring–Goldbach subbases
We are now in shape to prove Theorem 1.3. By the prime number theorem, $|\mathbb{P}^{k}\cap[1,x]|\sim kx^{1/k}/(\log x)$, which is a regularly varying function, so $\mathbb{P}^{k}$ satisfies condition (i) of Theorem 1.4. By Lemma 5.10, $\mathbb{P}^{k}$ also satisfies condition (ii), with $H_{\textnormal{(ii)}}\leq 2^{k-1}$.
For condition (iii), we claim that $H_{\textnormal{(iii)}}\leq 2^{k}+1$ for the functions considered in Theorem 1.3. Let $h\geq 2^{k}+1$, and $F(x)=x^{\kappa}\psi(x)$ be a regularly varying function with
$F(x)\leq(1+o(1))\frac{\Gamma(1/k)^{h}}{\Gamma(h/k)}\frac{x^{h/k-1}}{(\log x)^{h}}$
such that $F(x)(\log x)^{h}/x^{h/k-1}$ asymptotic to a non-increasing function. Note that the upper bound for $F$ comes from the total number of representations of $n$ as a sum of $h$ $k$-th powers of primes (cf. Hua [8, Theorem 11]). Define $f(x):=(xF(x))^{1/h}$, so that $f(x)\ll(1+o(1))\pi(x^{1/k})\asymp x^{1/k}/(\log x)$. For the $\eta=\eta(f,k,h)$ of Proposition 2.2, we have
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{P}^{k}\\ x_{1}+\cdots+x_{h}=n\\ \exists i\leavevmode\nobreak\ |\leavevmode\nobreak\ x_{i}<n^{\eta/2}\end{subarray}}\frac{f(x_{1})}{\pi(x_{1}^{1/k})}\cdots\frac{f(x_{h})}{\pi(x_{h}^{1/k})}$ $\leq\sum_{m\leq n^{\eta/2}}\frac{f(m)}{\pi(m^{1/k})}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h-1}\in\mathbb{P}^{k}\\ x_{1}+\cdots+x_{h-1}=n-m\end{subarray}}\frac{f(x_{1})}{\pi(x_{1}^{1/k})}\cdots\frac{f(x_{h-1})}{\pi(x_{h-1}^{1/k})}$
$\ll n^{\eta/2}\,n^{-\eta}\frac{f(n)^{h}}{n}=n^{-\eta/2}F(n),$
therefore
$\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{P}^{k}\\ x_{1}+\cdots+x_{h}=n\end{subarray}}\frac{f(x_{1})}{\pi(x_{1}^{1/k})}\cdots\frac{f(x_{h})}{\pi(x_{h}^{1/k})}=\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{P}^{k}\\ x_{1}+\cdots+x_{h}=n\\ \forall i,\,x_{i}\geq n^{\eta/2}\end{subarray}}\frac{f(x_{1})}{\pi(x_{1}^{1/k})}\cdots\frac{f(x_{h})}{\pi(x_{h}^{1/k})}+O(n^{-\eta/2}F(n))$
$\hskip 20.00003pt=(1+o(1))\frac{1}{k^{h}}\sum_{\begin{subarray}{c}x_{1},\ldots,x_{h}\in\mathbb{P}^{k}\\ x_{1}+\cdots+x_{h}=n\\ \forall i,\,x_{i}\geq n^{\eta/2}\end{subarray}}\frac{f(x_{1})}{x_{1}^{1/k}}\cdots\frac{f(x_{h})}{x_{h}^{1/k}}(\log x_{1}\cdots\log x_{h})+O(n^{-\eta/2}F(n))$ (6.1)
Choosing $F(x):=cx^{\kappa}$ with $10<\kappa<h/k-$ 1 and $c>0$, we get $f(x)=c^{1/h}x^{\omega}$ with $\omega=(1+\kappa)/h$, so $1/h<\omega<1/k$. Therefore, Theorem 1.2 together with (6.1) implies that for such $F$, $\mathbb{P}^{k}$ satisfies condition (iii).
Choose $F(x):=x^{\kappa}\psi(x)$ for some $10\leq\kappa\leq h/k-$ 1 and some measurable positive real function $\psi$ satisfying $\psi(x)\asymp_{\lambda}\psi(x^{\lambda})$ for every $\lambda>0$. Suppose that $\log n\ll F(n)\ll x^{h/k-1}/(\log n)^{h}$. We have $f(x)=x^{(1+\kappa)/h}\psi(x)^{1/h}$, where $1/h\leq(1+\kappa)/h\leq 1/k$. By uniform convergence,555[1, Theorem 2.0.8] applied to $g(x):=\psi(e^{x})$, so that $g(\lambda x)\asymp_{\lambda}g(x)$. it follows that $\psi(x_{i})\asymp\psi(n)$ for $n^{\eta/2}\leq x_{i}\leq n$, so Theorem 1.2 implies that the sum in (6.1) is $\asymp n^{\kappa}\psi(n)$ for $n\equiv h\leavevmode\nobreak\ (\mathrm{mod}\leavevmode\nobreak\ K(k))$. This fits into condition (iii) (with “$\mathfrak{S}_{B,h}(n)$” different from $\mathfrak{S}^{*}_{k,h}(n)$, but still $\asymp 1$).
Since $\max\{2H_{\textnormal{(ii)}}+1,H_{\textnormal{(iii)}}\}\leq 2^{k}+1$, the conclusion of Theorem 1.3 follows.
Acknowledgements
I thank Andrew Granville for his helpful advice during the preparation of this article. I also appreciate Cihan Sabuncu for insightful discussions that contributed to this work.